A Complete Bibliography of Publications in the Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches

sur les pêcheries du Canada (1950–1959)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

1 [134]. _{12} [47, 52, 68, 124]. × [140].

-Active [52].

0^{o} [295, 277].

15^{o} [295, 277]. 1950 [44]. 1955 [426].

Abundance [200, 117, 131, 303]. Acanthocephala [354]. Acclimation [179]. Acetone [251]. Acid [203, 260, 279, 291, 415, 1, 8, 98]. Acid-Soluble [8, 98]. Acids [432, 329]. Acipenser [195]. Actin [399]. Action [276, 288, 382, 383, 69]. Active [395, 52]. Activity [260, 279, 291, 413, 47, 374, 99, 433]. Adaptation [115]. Additional [442]. Additions [129]. Adjacent [239, 312]. Adjoining [197]. Adult [118, 122, 403]. aeglefinus [367, 235, 320]. Affected [187]. Affecting [421, 64, 359, 154]. after [16, 137]. Age [319, 87, 204, 117, 367, 372, 157, 317, 190]. Age-Composition [190]. Agent [217]. Agonid [49]. aix [49]. Akpatok [198]. alalunga [157, 158]. Alaska [221, 407]. Albacore [157, 158]. Alberni [208]. Albumin [189]. Algonquin [101]. along [192, 299, 324]. alpinus [253, 87]. Alterations [260, 279, 291]. Altered [290]. Ambient [395]. America [136, 148, 116, 151, 306]. American [222, 171, 102, 215, 88]. americanus [374, 88]. Amines [4, 27, 28]. Amino [432, 1]. Among [264, 420, 321]. Amphipod [143]. Amphipoda [17, 119, 164]. Anadromous [420, 53]. analoga [435]. Analysis [338, 428, 376]. Androgen [424]. Androgens [107]. Angling [36, 170]. Anisakis [185]. Annelids [221, 133]. Annual [104, 405, 422, 418, 141]. Anomaly [147]. Anomuran [435]. Antibiotics [287]. Appearance [320]. Appendicularians [128]. Applications [275]. Approach [90]. April [44]. Aquarium [237]. Archipelago [286]. Arctic [253, 345, 196, 143, 244, 87, 418, 414, 56, 42, 35]. Area [265, 200, 56, 367, 316, 284, 141, 436]. Areas [317, 113, 293]. Aricidea [239]. Arm [191]. Arthropods [236]. Artificial [159, 328]. Ascarid [324, 302]. Aspects [118, 402, 313, 77]. Assessment [241]. Associated [394]. Atlantic [324, 220, 397, 56, 371, 125, 192, 377, 113, 149, 229, 299, 241, 195, 381]. atromaculatus [111, 111]. August [426]. Aureomycin [380]. Average [405, 422, 444]. ayresii [306].

B [47, 134, 52, 68, 124]. B.C. [297, 319, 248, 147, 131]. Babine [162, 248]. baccatum [153, 154]. Bacteria [218, 51, 115, 216, 217, 69]. Bacterial [2, 21, 287, 67]. Bacto [216, 217]. Bacto-Oxgall [216, 217]. Baffin [87, 226]. Bank [317]. Banks [221, 267]. barbata [49]. barbatus [313]. Barkley [352, 402, 403]. Barnacles [70]. Basic [208]. Basicity [98]. Basin [176, 363, 308]. Bass [167]. Bathypelagic [132]. Bay [143, 311, 167, 87, 133, 202, 302, 104, 226, 296, 200, 46, 311, 201, 207, 133, 109, 202, 366, 72, 370, 309, 42, 141, 379]. Bear [66]. Bearded [313]. Beaufort [221]. Beds [152]. been [290]. Behaviour [23, 108, 227, 321, 172]. Belle [362, 114, 128, 17]. Below [346]. Benthos [54]. Between [135, 24, 443, 378, 383, 194]. Bibliography [408]. Bight [261]. Bilgy [254]. Biochemical [415, 398, 424]. Biological [251, 182, 438]. Biology [253, 201, 56, 371, 154]. Bladder [320]. Blennioid [434]. Blood [203, 260, 279, 291, 413]. Board [205, 206, 256, 300, 301, 411, 445, 446, 414, 361, 342]. Body [18, 398, 194]. Bonavista [220]. Boreo [56]. Boreo-Arctic [56]. Bottle [349, 358]. Bottles [307]. Bottom [257, 34, 126, 178, 243, 92]. Boundaries [225]. Bridge [409].

Brines [60]. British [142, 12, 63, 348, 400, 162, 351, 352, 247, 223, 250, 430, 427, 139, 270, 322, 232, 268, 269, 166, 49, 434, 185, 79, 435, 48, 405, 422, 403, 121, 62, 280, 34, 52, 318, 64, 410, 238, 73, 275, 86, 303, 443, 190, 273]. Brook [378, 53]. Brünnich’s [198]. Brunswick [145, 372, 439, 36, 31]. Bubbles [96]. Budget [275, 322]. Burbot [204]. Bute [223, 270]. Butter [33].

C [346]. Cabot [233]. Caeca [99]. caerulea [237]. Calanoida [201]. Calanus [45, 311, 210, 201, 109]. callarias [432, 377]. Canada [205, 300, 411, 445, 129, 221, 414, 361, 119, 342, 171, 381, 206, 256, 301, 446]. Canadian [345, 143, 419, 192, 299, 324, 259, 35, 286, 196, 156, 244, 418, 32, 186, 163, 236, 371, 423, 153]. Candling [327]. Cannery [308]. Cape [262]. Capelin [317, 316]. Car [277]. Carbon [344, 395]. Carbonate [413]. Cards [207]. Catch [161, 91, 101, 190]. Catchability [374]. Catches [319]. caudal [436]. Caught [422, 116, 416, 390, 397]. Cavity [145]. celata [145]. Central [70]. Certain [260, 279, 291, 287, 271, 77, 423, 171, 68, 359, 40, 101]. Cestode [162, 148]. Cetacea [259]. Chaetognatha [407]. Chaetognaths [186]. Change [178]. Changes [334, 18, 413, 241]. Channel [44, 369, 55, 208]. Char [253, 87, 140, 42]. Characteristics [132, 172, 266]. Characters [151]. Charles [262]. Charlotte [265, 36]. Chars [151]. Chatham [232]. Chemical [418, 241, 208, 381]. Chilled [276]. Chinook [390]. Chitin [98]. Chitosans [93]. Chloride [432, 18, 61]. Chlortetracycline [380]. Choice [246]. Cholesterol [398]. Chubb [182]. Chum [344, 18, 346, 422, 23, 227, 440, 94, 64, 177, 107, 359]. Ciliata [166]. Circulation [307]. Cirripede [196]. Cirripedia [70]. Ciscoes [38]. Clam [375, 33]. clarkii [443, 443]. Climatic [334, 230]. Climatological [19]. Clipping [10]. Close [14]. Clupea [185, 428, 318, 234, 379]. clupeaformis [391, 89]. Coast [221, 239, 220, 116, 119, 86, 303, 42, 185, 156, 39, 236, 437, 410, 73]. Coastal [440, 186, 318, 77, 266]. Cod [417, 397, 160, 310, 399, 4, 295, 277, 255, 47, 192, 376, 327, 377, 149, 229, 299, 396, 293, 124, 241, 386]. Codfish [432, 329, 61, 159]. Coho [18, 118, 332, 85, 23, 94, 10]. Collapse [214]. Collection [221, 82, 97, 375]. Colorimetric [3, 6]. Colour [86]. Columbia [185, 410, 73, 166, 142, 49, 434, 79, 12, 435, 63, 348, 48, 400, 162, 351, 405, 422, 352, 403, 121, 247, 62, 223, 250, 280, 430, 34, 52, 318, 64, 238, 275, 427, 86, 303, 443, 139, 270, 322, 190, 232, 268, 273, 269]. Commentary [408]. Comments [129]. Commercial [334, 348, 393, 281]. Common [285, 302]. Comparative [113, 53]. Comparison [319, 423, 273, 195]. Comparisons [351, 368]. Competition [228, 305, 80]. Components [223]. Composition [329, 383, 1, 190, 381]. Compounds [8]. Concentrates [124]. Concentrations [395]. Concept [431]. Concerning [442, 272]. Condensed [429]. Condition [215]. Conditions [19, 30, 327, 380]. Conductivity [409]. Considerations [309, 272, 281]. Constant [315]. Content [18, 421, 28, 424, 68]. Contrasting [134]. Contribution [46]. Control [228, 328, 375, 170]. Conversion [425]. Copepod [145]. Copepods [265, 201]. Coregonids [274]. Coregonine [66]. Coregonus [183, 391, 89]. Corrections [205, 300, 411, 445]. Corrosion [165]. Corynosoma [354]. Count [192]. County [36]. Course [315]. Cowichan [63]. Crater [182]. Creek [85]. Cristivomer [146, 140, 151]. Crop [184, 81]. Cruising [332, 10]. Crustacea [196, 143]. Crustacean [435]. Cultured [346]. Cultus [85, 54]. Cumberland [311, 133, 202]. Current [264, 55, 94, 105, 427]. Currents [44, 103, 144, 233, 243]. Cut [337, 339, 4]. Cutthroat [137, 283, 443, 266]. Cyamidae [119, 164]. Cycle [418, 135]. Cyclopoida [201]. Cyclopterichthys [48].

Data [91, 323, 190, 35]. Day [172]. Days [295]. decipiens [95, 324, 302]. Deck [209]. Decline [65]. Deep [226, 132, 369]. Definitive [95]. Deflect [96]. Denaturation [29, 7]. Density [18, 353]. Deodorization [20]. Deoxyribonucleic [415]. Depth [444]. Description [239, 82]. Deshayes [33]. Désodorisation [20]. Destructive [134]. Detection [327]. Deterioration [278]. Determination [376, 275]. Determinations [367]. Developing [344, 264]. Development [383]. Diagrams [169]. Diameter [194]. Diet [237, 134, 404]. Diet-Induced [237]. Differences [237, 443]. Different [392, 137]. Differentiation [217]. Digby [200, 141]. Digging [33]. Diluted [223]. Dimensions [257]. Dioxide [344, 395]. Dips [219]. Direct [264]. Direct-Current [264]. Director [343]. Discoloration [278, 363]. discovery [41]. Dispersion [34]. Dissolved [314]. Distinct [306]. Distinctness [267]. Distribution [37, 265, 400, 333, 58, 402, 248, 353, 148, 271, 250, 280, 52, 131, 157, 303, 316, 274, 407, 379, 128, 153, 433]. Diversions [280]. Dodecaceria [120]. Dogfish [437, 123]. Dominant [362]. Downstream [137]. Drainage [250, 441]. Drift [296, 349, 207, 307, 358]. Drill [326, 375]. Drop [115]. Drumming [320]. Drying [159, 328]. During [413, 276, 397, 415, 277, 398, 424, 426, 244]. Dwelling [283]. Dynamics [281].

East [221, 239, 220]. Easterly [267]. Eastern [143, 438, 196, 240, 419, 157, 158, 259, 151, 153]. Echinoderms [202]. Ecology [183]. Economic [90, 161]. Economics [281]. Edible [145]. Editors [341, 360]. Eel [171]. Effect [297, 432, 332, 287, 240, 30, 67, 180, 327, 10, 380, 130, 433, 224, 216, 425]. Effects [289, 314, 364, 402, 280, 39, 222]. Efficiency [327]. Effluent [289]. Effort [24, 91]. Egg [173]. Eggs [344, 314, 16, 152]. Electrical [264, 231, 263]. Electrophoretic [160]. elongatus [348]. Emerita [435]. Encountered [277]. Energy [309]. Enrichment [441]. Entomostracan [131]. Enumeration [115]. Environment [402]. Environmental [359]. Environments [355]. Enzyme [425, 99, 433]. Enzymes [21, 287]. Enzymic [188]. Eopsetta [298]. epizoum [150]. Equilibrium [169]. Erignathus [313]. Errors [388]. Erxleben [313]. Escapement [37]. Esox [245, 32]. Estimates [388, 348, 180]. Estimating [91, 340]. Estimation [251, 26, 304, 3, 27, 6, 190]. Estuary [439, 175]. Eumetopias [303]. Europe [306]. European [102]. Evaluation [428]. evides [434]. Evolution [321, 274]. Examination [185]. Exchanges [72]. Exercise [203]. Expeditions [45, 311, 109, 210]. Experimental [149]. Experiments [298, 96, 296, 26, 255, 338, 95, 423]. Exploitation [282]. Exploits [191]. Exposure [314]. Extractable [160]. Extraction [251, 387, 310, 5]. Extractives [387]. Extracts [384, 188]. extrusion [436].

F [295, 277, 295, 277]. Fabricius [215]. Factor [331, 15, 428]. Factors [265, 421, 161, 247, 271, 83, 268, 359]. Family [434]. Farm [441]. Fasting [67]. Fat [387, 398, 428]. Fate [432]. Fatigue [404, 416, 390]. Fatness [370]. Fats [251]. Fatty [329]. Fauna [129, 394, 257, 92]. Features [286, 426, 175]. Fecundity [89, 318, 252]. Fed [237]. Feeding [326, 87, 135, 377, 198]. female [436]. Fertility [16]. Fertilization [170]. fewkesi [120]. Field [414, 409]. Fillets [337, 339, 421, 4, 213, 242, 327, 299, 396, 293, 241, 433]. Fin [10]. finmarchicus [201]. First [434, 48, 102]. Fish [49, 132, 434, 395, 330, 331, 184, 14, 50, 209, 219, 224, 276, 288, 337, 339, 382, 383, 421, 397, 401, 251, 387, 26, 310, 386, 399, 29, 51, 216, 217, 4, 5, 27, 28, 385, 384, 333, 188, 394, 134, 257, 338, 398, 111, 228, 328, 280, 429, 67, 425, 254, 22, 423, 6, 7, 112, 8, 11, 235, 320, 380]. Fisheries [205, 206, 256, 300, 301, 411, 445, 446, 334, 282, 414, 230, 281, 361, 342]. Fishery [169, 167, 90, 391, 371, 214, 181, 1, 68, 246]. Fishes [142, 420, 79, 203, 260, 279, 291, 392, 15, 46, 110, 66, 419, 250, 34, 439, 97, 290, 410, 431, 293]. Fishing [264, 393, 24, 290, 376, 340, 113, 284]. Fjord [249]. Flatfish [165, 91]. Flavobacterium [50]. Flesh [417, 62, 336]. Flora [67]. Flounder [169, 39, 113, 154]. Flounders [153]. Flow [362, 267]. Fluctuating [355]. Fluctuations [200, 161, 117, 230, 78]. Fluviarium [335]. fluviatilis [306]. Folliculinids [166]. Following [203, 260, 279, 291, 413, 264, 168]. fontinalis [179, 140, 378, 78, 252, 53]. Food [204, 32, 101, 198]. Foreword [341, 360]. Four [394]. Foxe [363]. Fractions [189, 5]. fragi [383]. Francis [261]. Fraser [308]. Free [344]. Freezing [14, 7]. French [20]. Frequency [33]. Fresh [18, 5, 72, 275, 171, 378]. Freshly [4]. Freshwater [142, 79, 203, 401, 228, 439, 380, 53]. Frobisher [311, 133, 202, 87, 109]. Frozen [278, 213, 242, 295, 277, 211]. Fruity [382]. Fry [18, 227, 240, 94, 176, 177, 10]. Fuca [103, 144]. Fundy [104, 296, 200, 366, 72, 370, 309, 302, 141, 379]. Further [295, 43].

Gadus [432, 377]. gairdneri [194, 260, 413]. Gastropoda [173]. Gay [296]. Gear [264]. General [114]. Genetic [290]. Genus [57, 321, 148]. Geographical [148]. Georgia [348, 268, 269, 132, 271, 243, 358]. Georgian [207]. Giant [200, 364, 141]. giganteus [33]. Gilbert [434]. Gill [24, 406]. Gills [224]. Glands [106]. Glycogen [404]. Gmelin [200, 157, 141]. Goldfish [10]. Gonads [158]. Graded [337]. Grades [338]. Grading [337, 339, 338, 423, 339]. Grand [55, 267, 317]. Gravel [193, 357]. Great [197, 58, 89, 66, 146, 59, 9, 22, 92, 212]. Greater [375]. groenlandica [20]. Groundwater [357]. Growth [392, 15, 245, 87, 32, 106, 38, 89, 146, 34, 83, 372, 313, 431, 157, 317, 80, 140, 141, 235, 88]. Gulf [377, 396, 294, 407, 379, 195]. Gulls [285]. Günther [306]. Gutted [219, 325].

Habitable [187]. Habitat [394]. Habits [326, 87, 198]. Haddock [417, 174, 397, 255, 135, 368, 367, 376, 235, 325, 320]. Halibut [371]. Halophiles [60]. Halophilic [51, 115, 216, 217]. Harbour [261, 95]. harengus [234, 379]. Harpacticoid [145]. Hatchery [305, 404]. Head [296, 231, 263]. Hearing [111]. Heat [275, 322]. Hebron [249]. Hecate [297, 247, 105]. Helminth [163]. Helminths [236]. Heming [100, 312]. Hemoglobin [203]. heros [375]. Herring [185, 352, 402, 403, 65, 62, 370, 428, 373, 285, 318, 406, 190, 234, 379]. heterotricha [166]. Heterotrophic [218]. Hiemal [420]. High [364]. Hippoglossoides [215]. Hippoglossus [371]. Histological [158]. History [71, 120]. Hold [209, 219]. Homarus [374, 88]. Home [283]. Homes [137]. Homing [430]. Hormone [106]. Host [95]. Hosts [148]. Hudson [311, 133, 202, 42, 196]. Huile [20]. Humber [191]. Humpback [86]. Huron [167]. Hybrid [140]. Hydrogen [287]. Hydrolysates [188]. Hydrolysis [385]. Hydroxylamine [218, 69]. Hyperactivity [331].

Ice [363, 219]. Identification [382]. Identity [229]. Igloolik [418]. II [279, 217, 263, 1, 6, 154]. III [291, 7]. illecebrosus [284]. Illex [284]. Immersed [61]. Importance [50]. Impounded [176]. Improve [170]. Improved [51]. Incidence [299, 324, 302, 396, 308]. Incipient [276]. including [119]. Inconnu [197]. Incursion [75]. Index [206, 256, 301, 446, 351]. Indicated [132, 307]. indicator [438]. Indicators [2]. Indices [238]. Induced [237]. Infection [162, 149]. Influence [2, 20, 295, 277, 20]. Influencing [265, 83, 268, 247]. ingens [41]. Injury [168]. Inlet [430, 427, 270, 350, 319, 208]. Inlets [223, 275]. Inner [366]. Inshore [297]. Interaction [401]. Intermittent [277]. Interspecific [228]. Interval [24]. Intestine [67]. Introductions [290]. Intrusion [308]. Invertebrates [28, 52]. Investigation [231, 263]. Investigations [414]. Iodide [310, 399]. Ionic [160, 223]. Island [221, 239, 87, 109, 43, 322, 198]. Islands [265]. Isle [17, 362, 114, 128]. Isopleth [169]. IV [29, 4, 425].

Japan [97]. Japanese [326]. japonica [326]. Jellied [215]. John [139]. Johnson [41]. jordani [298]. Journal [206, 256, 301, 446]. Juan [103, 144]. jubata [303]. Juvenile [108, 292, 321, 315, 352, 402, 403, 354].

Kamloops [260, 413, 194]. Keeping [437, 380]. Kerr [31]. keta [18, 107]. Kinds [119]. kisutch [18, 130]. Knight [223, 427]. Known [122]. Kokanee [442, 273]. Kraft [289].

L. [253, 367, 371, 377, 294, 235, 336, 320, 234, 379]. Laboratory [255]. Labrador [253, 262, 261, 77, 249]. Lack [179]. Lactate [413]. Lactic [203, 260, 279, 291]. Lake [12, 63, 400, 162, 167, 187, 257, 391, 353, 66, 146, 240, 100, 312, 430, 101, 131, 139, 31, 273, 279, 13, 85, 197, 204, 393, 391, 38, 89, 146, 117, 9, 22, 92, 212, 54]. Lakelse [13, 12, 131]. Lakes [184, 333, 248, 312, 101, 444, 238, 81, 36, 170, 258, 252, 58, 59]. Lampetra [306]. Lamprey [231, 263, 25, 306]. Landed [397]. Landings [325]. Lands [441]. Large [65, 273]. Larvae [185, 172, 285, 436]. Larval [149, 229, 299, 379]. late [436]. late-stage [436]. latipes [408]. Laurentian [369]. Lawrence [74, 377, 396, 294]. Learning [315]. lemon [247]. Length [237, 194, 320]. Lengths [235]. Lesser [258]. LeSueur [284]. Lethal [79, 331]. leucichthys [197, 38]. Leuciscus [335]. Level [260, 279, 291, 282, 257, 147]. Levels [314, 203]. lewisi [443]. Liberated [329]. Lice [119, 164]. Life [71, 120]. Life-History [71, 120]. Lifts [24]. Light [292]. Lighting [327]. Lightly [159]. Lights [96]. like [288]. Limnobiology [63]. Limnological [312, 182]. Limnology [12, 9, 36]. Ling [102]. Lingcod [348, 433]. Linnaeus [306]. Lion [303]. Lipase [433]. Lipid [385]. Lipoxidase [62]. List [45, 109, 210, 414, 116]. Liver [47, 188, 124]. Lobster [278, 222, 374, 285, 88]. Lobsters [40]. Lockeport [367, 396]. Lockington [298]. lomvia [198, 198]. Loon [430]. Lota [204]. Loudoun [44]. Low [314, 160, 179, 433]. Lower [302]. Lump [48]. Lump-Sucker [48]. Lunatia [375]. lyra [438].

Mackenzie [250]. mackenziei [197]. Magdalen [30, 74]. magellanicus [200, 141]. Maillard [11]. Maine [379]. Mainland [142, 192, 299, 324]. Major [403, 223]. Mallotus [317, 316]. Mammalian [188]. Mammals [116]. Management [281]. Manan [55]. Manitoba [100, 312]. Mantle [145]. Manual [375]. Marine [334, 28, 230, 52, 439, 281, 95, 149]. marinus [231, 263, 294, 336, 436]. marinus- [436]. Maritime [171]. Mark [357]. Marking [304]. Marks [25]. Maskinonge [245, 32]. masquinongy [245, 32]. Mass [296]. Masses [297, 173]. Material [47]. Materials [251, 68]. Mathematical [246]. Matter [258]. Maturity [319, 204, 89, 146, 320]. Maximizing [246]. Maximum [444, 355]. Mean [147]. Measurement [174, 323]. Measurements [55, 427, 208]. Measuring [193, 357]. Meat [278]. Medaka [408]. Medium [51]. Melanogrammus [367, 235, 320]. mentella [436]. mentella-type [436]. Mesh [406]. Metacercaria [153]. Metal [165]. Metals [40]. Method [115, 340, 180]. Methods [82, 323]. Mexico [195]. Micronereis [71]. Midsummer [333]. Migrating [96, 118, 227, 130]. Migration [415, 87, 23, 398, 424, 177]. Migrations [122, 373, 123]. Migratory [321]. Mill [289]. Mineralization [84]. Minimum [348]. minutus [201]. Miramichi [175, 372, 439]. Mitchill [245, 32, 111, 372, 378, 78, 53]. Mixed [355]. Mixing [127]. Mixtures [251]. Models [246]. Molecular [258]. Monstrilloida [201]. Months [397]. mordax [372]. Morphological [443]. Morphology [164]. Morphometric [273]. Morphometry [66]. Mortalities [264, 376, 340]. Mortality [388, 348, 245, 89, 146, 305, 416, 390, 190]. Movement [176]. Movements [245, 400, 207, 65, 137, 378]. Müller [317, 316]. Multibasin [353]. Murre [198]. Muscle [432, 276, 288, 382, 383, 387, 310, 386, 399, 29, 4, 5, 27, 28, 385, 384, 61, 6, 7, 8]. Muscles [320]. Muscular [260, 279, 291, 413, 390]. Mussel [145]. Musty [288]. Myosin [310, 7].

n [71, 120, 145]. N.S. [367]. N.W.T. [66]. namaycush [279, 146, 140]. nanaimoensis [71]. Narrows [147]. Naticidae [173]. Natural [47, 376, 340, 80, 170, 190]. Nature [258]. Neighbouring [293]. Nematoda [185, 229]. Nematode [162, 95, 149]. Nematodes [299, 396, 293]. nerka [291, 122, 155, 108, 248, 353, 442, 273]. Nervous [70]. Net [44, 81, 212]. Nets [24, 406]. Newfoundland [220, 317, 316, 284, 123, 293, 325, 436, 234]. Nicoll [153, 154]. Night [172]. Nilkitkwa [248]. Nitrite [218, 219, 421, 241]. Nitrite-Treated [241]. nom [120]. Non [296]. Non-tidal [296]. North [136, 116, 151, 306, 148, 102]. Northeast [426, 349, 230]. Northeastern [97]. Northern [434, 303, 253, 221, 311]. Northwest [221]. Note [344, 218, 189, 3, 399]. Notes [71, 120, 239, 148, 34, 164, 177]. Nova [376, 396, 187, 225].

O. [18, 130]. Obituary [199, 412]. Observations [347, 15, 418, 100, 176, 105, 182, 442, 358]. Obtained [388]. Occlusion [130]. Occurrence [354, 379]. Ocean [323, 426, 349, 56, 438]. Oceanic [97]. Oceanographic [286, 426, 311, 418, 271, 225, 35]. Oceanography [114, 77, 249, 270, 232, 269]. Oceans [258]. Ocinebra [326]. Odour [110]. Odours [276, 288, 382, 383]. Off [74, 296, 116, 76, 267, 97, 410, 376]. Offshore [132, 156, 316]. Oil [20]. old [389]. Older [129]. Olfactory [118, 130]. omiscomaycus [100]. Oncorhynchus [18, 291, 57, 122, 155, 108, 321, 248, 353, 354, 416, 390, 107, 442, 273]. Oniony [382]. Ontario [394, 240]. Ophiodon [348]. Optimum [282, 90]. Orange [417]. Orange-red [417]. Organ [168]. Organic [258]. Organisms [382, 276]. Orientation [110]. Origin [226, 267]. Oryzias [408]. Osmerus [372]. Osmophilism [150]. Ostracoda [43]. other [293]. Otolith [368, 367, 235]. Outlet [430]. Ova [264]. Oxgall [216, 217]. Oxide [2, 21, 27, 28]. Oxygen [314, 395, 222, 179, 152]. oxyrhynchus [195]. Oyster [326, 172]. Oysters [211].

Pacific [426, 349, 156, 236, 289, 314, 57, 347, 108, 292, 315, 230, 438, 176, 39, 119, 163, 354, 428, 437, 318, 416, 157, 158]. Pallas [48]. pallasi [185, 318]. pallasii [428]. Pallasina [49]. Parasites [142, 116, 163, 327]. Parasitic [236, 302]. Parasitism [38]. Park [101]. Parophrys [247, 39]. Particular [19, 319, 47, 108, 323, 20]. Particularly [251]. Particulier [20].

Passamaquoddy [72, 373]. Pattern [86]. Patterns [290]. Paul [400]. Pelagic [17, 265]. Penetrating [323]. Perception [118, 110]. Perch [100]. Percopsis [100]. Period [397]. Periodicity [272]. Periods [295, 277, 137]. Permanence [283]. Petersen [180]. Petersen-type [180]. Petrale [298]. Petromyzon [231, 263]. Petromyzontidae [306]. pH [413]. Phoca [20, 95]. Phocoena [302]. Phoque [20]. Phospholipids [329]. Phosphorous [8]. Phosphorus [441]. Photomechanical [347]. Photosynthetic [323]. Phototactic [168]. Physical [175, 12, 311, 418, 9, 270, 269]. Phytoplankton [83]. Pigmentary [168]. pigmentation [436]. Pineal [168]. Pink [389, 405, 23, 227, 440, 64, 177, 359]. Pinnipedia [236]. Pituitary [106]. Placopecten [200, 141]. Plaice [213, 215]. Plankton [84, 82, 271, 131, 81, 212]. Planktonic [201]. Planning [26]. Planted [125]. Plantings [240]. Plate [115]. platessoides [215]. Plecoptera [136]. Plectobranchus [434]. Poisoning [187]. Pollution [330, 191]. Polychaeta [239, 345]. Polychaete [41, 129, 120]. Polychaetous [221, 133]. Population [245, 304, 187, 352, 91, 228, 180, 281]. Populations [388, 26, 403, 192, 64, 80, 190, 234, 53]. Porpoise [302]. Porrocaecum [95, 229, 324, 302]. Port [139]. Portable [409]. Possible [280, 285]. Post [160, 384]. Post-Rigor [160, 384]. Potassium [310, 399]. Potato [288]. Potato-like [288]. Potential [87]. Potentials [231, 263]. pre [436]. pre-extrusion [436]. Predator [400, 401, 170]. Predators [285]. Predict [169]. Predicting [246]. Preference [335]. Preferred [333]. Preliminary [96, 326]. Preparation [329, 399, 93, 429, 124]. Presence [107]. Present [432]. Preservation [219]. Prevailing [297]. Prey [400, 401]. Principles [64]. Procedure [304]. Processing [255]. Produced [276, 231, 263]. Producing [382]. Production [19, 218, 287, 440, 240, 356, 359]. Productivity [238, 33, 323, 208]. Products [251, 1, 11]. Progress [349]. Properties [20, 429, 268]. Propriétés [20]. Prosopium [183]. Protein [189, 5, 384, 47, 398, 6]. Proteins [160, 310, 386, 399, 29, 5, 385, 384, 6, 7]. Proteolysis [432]. Proteolytic [99]. Province [182]. Provinces [171, 381]. Proximate [428]. Pseudocalanus [201]. Pseudomonas [276, 288, 382, 383]. publications [205, 300, 411, 445]. Pyloric [99].

Qualitative [271]. Quality [174, 337, 339, 397, 295, 277, 338, 437, 423, 325, 380]. Quantitative [271]. Quantity [325]. Quebec [182, 252]. Queen [265]. Quelques [20]. Quoddy [65].

Rôle [305]. Races [420, 273]. Racial [316]. Radiation [323]. Rainbow [400, 430, 404]. Rapid [251, 315]. Rate [224, 141, 235, 190, 88]. Rates [348]. Ratio [387]. Ratios [155, 223]. Re [41]. Re-discovery [41]. Reaction [11]. Reactions [292]. Recapture [304]. Recently [339]. Reception [112]. Record [434, 435, 48, 102]. Recorded [116]. Records [129, 221, 244, 419, 410, 376]. Recovery [214]. Recruitment [180, 138, 190].

Red [60, 51, 115, 216, 217, 336, 417]. Redescription [306]. Redfish [294, 99, 336, 436]. Redox [2]. Redside [400]. Reduction [2, 21]. Reference [19, 201, 319, 47, 108, 111, 181, 179, 323, 258, 154]. Refrigerator [277]. Regarding [338]. Region [297, 17, 196, 231, 263, 182, 373, 43, 35]. Relating [15]. Relation [395, 348, 122, 257, 23, 135, 38, 271, 374, 33, 281, 194, 308, 128, 241, 383]. Relations [127]. Relationship [237, 403, 24, 235]. Relative [140, 325, 40]. Relatively [89, 146]. Releases [349]. Reminiscences [343]. Removal [224]. Removing [180]. Report [349]. Reproduction [313, 377]. Reproductive [87]. Requirements [323]. Research [205, 206, 256, 300, 301, 411, 445, 446, 414, 361, 342]. Reserves [404]. Residual [442]. Residues [124]. Resistance [346, 165, 404, 179]. Resolution [109]. Resources [90]. Respect [423]. Respiration [395]. Response [344]. Responses [347, 168, 94, 112]. Responsible [276]. Resultant [356]. Results [298, 296, 311, 338]. Retention [137]. Retina [347]. Return [351]. Returns [122, 117]. Review [429, 323, 359]. Ribonucleic [415]. Rigor [160, 384]. Rings [141]. River [319, 223, 372, 439, 19, 37, 183, 181, 308, 191]. River-Diluted [223]. Rivers [350, 319]. Roaches [335]. Rockfish [121]. Rosefish [242]. Round [325]. rubrivinctus [121]. Run [74]. Run-Off [74]. rutilus [335].

sa [20]. Sagitta [438]. Salinities [392, 74]. Salinity [104, 244, 222, 127, 73]. Salmo [260, 413, 443, 194]. Salmon [289, 344, 314, 389, 18, 291, 220, 262, 261, 57, 96, 118, 332, 346, 347, 415, 85, 122, 155, 350, 264, 351, 319, 405, 422, 23, 108, 168, 227, 292, 321, 315, 440, 365, 398, 424, 106, 248, 353, 94, 176, 354, 125, 181, 64, 177, 416, 390, 193, 107, 10, 357, 152, 359, 130]. Salmonoid [16, 260, 279, 291, 15]. Salt [29, 51, 216, 217, 165, 72, 328, 378]. Salted [329, 216, 217, 159]. Salvelinus [253, 279, 87, 179, 140, 378, 78, 151, 252, 53]. Sampler [82]. Sanctuary [245]. Sardines [237]. Sardinops [237]. Saxidomus [33]. Scale [368, 194, 266]. Scallop [200, 364, 141]. Schools [315]. Scotia [187, 225, 376, 396]. Scotian [75, 76, 126, 127, 178, 267, 307]. Scottish [423]. Sea [18, 174, 209, 219, 224, 397, 223, 231, 263, 147, 116, 254, 303, 308, 195, 381, 221]. Seal [20, 313, 95]. Seals [324]. Season [320]. Seasonal [73, 78]. Seasons [393]. Seaward [122, 23, 177]. Sebastes [294, 336]. Sebastodes [121]. Seepage [193, 357]. Selected [78]. Selection [406]. Selective [217, 290]. Semotilus [111]. Seven [391]. Seven-year [391]. Several [351, 164]. Severe [413]. Sewage [258]. Sex [155, 320]. Sexual [135, 320]. Seymour [147]. Shallows [187, 30, 74]. Sharpness [225]. Shelf [75, 76, 126, 127, 178, 267, 307]. Shiner [400]. Shipworms [308]. Shock [264]. Short [295, 277]. shown [192]. Shubenacadie [365]. Significance [19]. Silver [130]. Six [393]. Size [348, 304, 283, 64, 356]. Skeena [356, 19, 37, 183, 319, 181]. Skeletal [8]. Slave [197, 89, 146, 9, 22, 92, 212]. Slope [75, 76, 267]. Small [169, 370, 176, 214, 396]. Smaller [259].

Smallmouth [167]. Smelt [372, 149]. Smolt [122]. Smolts [162, 168]. Smooth [48]. Sockeye [13, 19, 37, 332, 346, 415, 162, 122, 155, 350, 319, 108, 168, 398, 248, 353, 181, 442, 356, 291]. Sodium [432, 61]. Soekeye [424]. Soils [258]. Solar [216, 217, 323]. Sole [298, 247]. Soluble [8, 98]. Solubles [429, 425, 387]. Solution [344, 246]. Solutions [61]. Solvent [387]. Some [289, 314, 334, 426, 129, 79, 175, 118, 347, 265, 218, 15, 20, 244, 402, 223, 176, 163, 313, 443, 272, 281, 71, 221, 203, 34, 52]. sotoi [195]. Sound [311, 352, 232, 133, 202, 402, 403]. South [225, 267, 221, 167]. Southern [192, 299, 324]. Southwestern [377, 396]. sp [71, 145]. Spatfall [172]. Spawning [298, 415, 204, 398, 424, 430, 316, 193, 272, 356, 357, 152]. Special [201, 111, 181, 179, 154]. Species [239, 394, 134, 119, 164, 324, 306, 438]. Specific [229]. Speckled [179, 78, 252]. Spectrum [98]. Speed [332, 10]. Sperm [16]. Spike [231, 263]. Spiny [123]. Spoilage [50, 174, 209, 219, 224, 276, 397, 254, 241, 4]. Sporendonema [150]. spp [354]. Spring [96, 118, 74]. Squid [284]. St [377, 396, 294]. St. [74]. Stabilité [20]. Stability [20]. Stage [320, 436]. Stages [154]. Standard [398]. Standards [174]. Standing [184, 81]. Standpipe [357]. Starks [49]. Station [45, 109, 210, 414]. Stations [109]. Statistical [185, 338]. Status [139]. Steelhead [266]. Stenodus [197]. Stephanostomum [153, 154]. Steveston [308]. Stichaeidae [434]. Stickwater [425]. Stimpson [435]. Stock [298, 138, 356]. Stocking [187]. Stocks [403, 355, 351]. Stomachs [324]. Stoneflies [136]. Storage [16, 14, 213, 242, 295, 277, 433]. Stored [295, 277]. Storms [30]. Strait [297, 133, 202, 247, 362, 132, 17, 196, 348, 311, 103, 144, 114, 271, 233, 105, 243, 268, 128, 269, 358]. Stream [440, 125, 283]. Stream-Dwelling [283]. Streams [333, 394, 430]. Strength [160]. Structural [384]. Structure [347]. Studied [335]. Studies [415, 160, 384, 352, 398, 424, 117, 312, 425, 125, 192, 158, 316, 377, 22, 153, 154]. Study [19, 326, 46, 216, 217, 295, 351, 393, 391, 62, 113, 99, 53]. Studying [296]. Sturgeon [189, 272, 195]. subarctic [438]. Subspecies [443, 195]. Substances [52]. Substrate [383]. Sucker [48]. Sulphide [287]. Summer [74]. Supply [152]. Surface [103, 56, 59, 73, 243]. Surrounding [231, 263]. Survey [142, 191]. Survival [13, 60, 364, 351, 440, 222, 247]. Susceptibilities [134]. Sustained [355]. Sweltzer [85]. Swim [320]. Swim-Bladder [320]. Synonymy [164]. System [353, 309, 70].

Tag [376]. Tagged [220, 123]. Tagging [298, 262, 261, 165, 117, 91, 39]. Taxonomic [151]. Taxonomy [38, 250, 164]. Technique [296, 335]. Temperate [333]. Temperature [104, 334, 57, 332, 364, 20, 244, 333, 58, 222, 374, 127, 73, 78, 112, 433, 20]. Temperature-Salinity [127]. Temperatures [79, 346, 14, 126, 178, 59]. Temporary [314]. Temps [20]. Terns [285]. Terrestrial [258]. Territory [283]. Tester [428]. Testes [424]. their [19, 333, 23, 137]. Theoretical [348]. Theory [80]. Thermocouple [328]. Three [389, 119, 113, 324, 273]. Three-year-old [389]. throughout [353]. Thunnus [157, 158]. Tidal [103, 144, 105, 296]. Time [20]. Tisbe [145]. Together [221, 109]. Tolerance [57, 179]. Toleration [330]. Toxic [330]. Toxicity [40]. Transfer [61]. Transferred [18]. Transport [44, 233, 275]. Transportation [277]. Travelling [315]. Trawler [397]. Trawler-caught [397]. Trawlers [174]. Treated [421, 241]. Treatment [425, 209]. Trematode [153, 154]. Trends [230]. Trevor [208]. Triaenophorus [38, 148]. Trial [375]. Trimethylamine [2, 21, 287, 3, 338, 337, 27, 28]. Trinitrophenol [335]. Triple [322]. Troll [416, 390]. Troll-Caught [416, 390]. Trophic [257, 139]. Tropical [380]. Tropomyosin [386]. Trout [260, 279, 413, 400, 264, 187, 146, 100, 430, 101, 137, 283, 305, 404, 443, 179, 36, 170, 194, 378, 78, 266, 252, 53]. Trout-Perch [100]. Trypanosyllis [41]. tshawytscha [390]. Two [134, 312]. type [180, 436]. Types [165].

Ultracentrifuge [384]. Ultraviolet [98]. Uncommon [419]. Unexploited [89, 146]. Ungava [46, 143, 311, 201, 133, 202, 182]. Unidentified [15]. Unit [414]. Unusual [419]. upon [39]. Upper [79]. Upstream [137]. Uria [198]. Use [409, 91]. Used [165]. Using [96]. Utilization [90].

V [27, 398]. Validity [367]. Value [14, 219]. Values [337, 338]. Vancouver [239, 43]. Variation [422, 257, 73, 299, 320]. Variations [104, 351, 405, 436, 397]. Various [165, 327]. ventricosus [48]. Vernal [420]. Vertebral [192]. Vertical [58]. Vessel [219]. Vessels [209, 219, 224, 397, 254]. vetulus [247, 39]. VI [415, 28, 357]. Vicinity [322, 381]. VIII [424]. villosus [317, 316]. Virtual [388]. Viscera [255]. Viscosity [93, 425]. Vitamin [47, 134, 52, 68, 124]. vitulina [95]. Volume [256, 301, 446, 320, 206].

Walbaum [100, 107]. Water [226, 297, 18, 60, 251, 387, 165, 61, 75, 187, 398, 223, 231, 263, 30, 280, 176, 76, 267, 275, 378, 322, 308, 268, 381]. Water-solubles [387]. Waters [132, 434, 239, 220, 48, 156, 197, 422, 121, 72, 369, 186, 318, 410, 77, 73, 171, 441, 102, 123, 234, 438, 32, 419, 259]. Weeks [295]. Weight [237, 422]. Weights [405, 235]. West [116, 42]. Western [345, 136, 186, 376, 294, 35, 306]. Whale [119, 164, 425]. Whale-Lice [119, 164]. Whales [163, 25, 86]. Whitefish [391, 89, 117, 240, 214, 31]. Whitefishes [183]. Wild [252]. Winds [297]. Winnipeg [204, 393, 391, 38, 117]. Winter [75, 393, 101, 154]. Wire [165]. Within [378, 273]. Worm [41].

XII [206].

year [389, 391]. Yearling [413, 266]. Yellow [278]. Yellowtail [113]. Yield [348, 257, 230, 246]. Yields [334, 169, 355]. Young [289, 57, 96, 332, 346, 15, 248, 353, 179].

Zone [56]. Zooplankton [418, 366, 56].

References

Ney:1950:AAC

[1] Phyllis W. Ney, Catherine P. Deas, and H. L. A. Tarr. Amino acid composition of fishery products (II). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7d(10):563–566, October 1950. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f47-048.

Castell:1950:ITO

[2] C. H. Castell. The influence of trimethylamine oxide on the bacterial reduction of redox indicators. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7d(10):567–575, October 1950. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f47-049.

Dyer:1950:NCE

[3] W. J. Dyer. Note on the colorimetric estimation of trimethylamine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7d(10):576–579, October 1950. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f47-050.

Dyer:1950:AFM

[4] W. J. Dyer and F. E. Dyer. Amines in fish muscle: IV. Spoilage in freshly cut cod fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7d(10):580–584, October 1950. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-051.

Dyer:1950:PFM

[5] W. J. Dyer, H. V. French, and J. M. Snow. Proteins in fish muscle: I. Extraction of protein fractions in fresh fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7d(10):585–593, October 1950. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f47-052.

Snow:1950:PFMa

[6] J. M. Snow. Proteins in fish muscle: II. Colorimetric estimation of fish muscle protein. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7d(10): 594–598, October 1950. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f47-053.

Snow:1950:PFMb

[7] J. M. Snow. Proteins in fish muscle: III. Denaturation of myosin by freezing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7d(10):599–607, October 1950. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-054.

Tarr:1950:ASP

[8] H. L. A. Tarr. The acid-soluble phosphorous compounds of fish skeletal muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7d(10):608–612, October 1950. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-055.

Rawson:1950:PLG

[9] D. S. Rawson. The physical limnology of Great Slave Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8a(1):3–66, January 1950. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f50-001.

Radcliffe:1950:EFC

[10] Roland W. Radcliffe. The effect of fin–clipping on the cruising speed of goldfish and Coho salmon fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8a(2):67–73, February 1950. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f50-002.

Tarr:1950:MRF

[11] H. L. A. Tarr. The Maillard reaction in fish products. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8a(2):74–81, February 1950. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f50-003.

Brett:1950:PLL

[12] J. R. Brett. The physical limnology of Lakelse Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8a(2):82–102, February 1950.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f50-004.

Brett:1950:LLS

[13] J. R. Brett and J. A. McConnell. Lakelse Lake sockeye survival. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8a(2):103–110, February 1950. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f50-005.

Castell:1950:VTC

[14] C. H. Castell and W. A. MacCallum. The value of temperatures close to freezing on the storage of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8a(2):111–116, February 1950. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f50-006.

Cooke:1950:SOR

[15] Norman E. Cooke and R. V. Tomlinson. Some observations relating to an unidentified growth factor for young salmonoid fishes. I. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8a(2):117–123, February 1950. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f50-007.

Barrett:1951:FSE

[16] I. Barrett. Fertility of salmonoid eggs and sperm after storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(3):125–133, March 1951. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f50-008.

Bousfield:1951:PAB

[17] E. L. Bousfield. Pelagic Amphipoda of the Belle Isle Strait region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(3):134–163, March 1951. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f50-009.

Black:1951:CBC

[18] Virginia Safford Black. Changes in body chloride, density, and water content of chum (Oncorhynchus keta) and Coho (O. kisutch) salmon fry when transferred from fresh water to sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(3):164–177, March 1951. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f50-010.

Brett:1951:SSR

[19] J. R. Brett. A study of the Skeena River climatological conditions with particular reference to their significance in sockeye production. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(3):178–187, March 1951. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f50-011.

Dugal:1951:ITD

[20] L.-C. Dugal and A. Cardin. Influence de la température et du temps de désodorisation sur quelques propriétés de l’huile de phoque (Phoca groenlandica), et en particulier sur sa stabilité. (French) [Influence of temperature and deodorization time on some properties of seal oil (Phoca groenlandica), and in particular on its stability]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(4):189–194, April 1951. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f50-012.

Castell:1951:RTO

[21] C. H. Castell and J. Murray Snow. Reduction of trimethylamine oxide by bacterial enzymes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(4): 195–206, April 1951. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f50-013.

Rawson:1951:SFG

[22] D. S. Rawson. Studies of the fish of Great Slave Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(4):207–240, April 1951. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f50-014.

Hoar:1951:BCP

[23] William S. Hoar. The behaviour of chum, pink and Coho salmon in relation to their seaward migration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(4):241–263, April 1951. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f50-015.

Kennedy:1951:RFE

[24] W. A. Kennedy. The relationship of fishing effort by gill nets to the interval between lifts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(4): 264–274, April 1951. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f50-016.

Pike:1951:LMW

[25] Gordon C. Pike. Lamprey marks on whales. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(4):275–280, April 1951. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f50-017.

DeLury:1951:PEE

[26] D. B. DeLury. On the planning of experiments for the estimation of fish populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8b(4): 281–307, April 1951. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f50-018.

Dyer:1952:AFMa

[27] W. J. Dyer, F. E. Dyek, and J. M. Snow. Amines in fish muscle. V. Trimethylamine oxide estimation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(5):309–313, May 1952. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f50-019.

Dyer:1952:AFMb

[28] W. J. Dyer. Amines in fish muscle. VI. Trimethylamine oxide content of fish and marine invertebrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(5):314–324, May 1952. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f50-020.

Duerr:1952:PFM

[29] J. D. Duerr and W. J. Dyer. Proteins in fish muscle. IV. Denaturation by salt. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(5):325– 331, May 1952. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f50-021.

Lauzier:1952:ESW

[30] L. Lauzier. Effect of storms on the water conditions in the Magdalen Shallows. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(5):332– 339, May 1952. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f50-022.

Smith:1952:LWK

[31] M. W. Smith. The lake whitefish in Kerr Lake, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(5):340–346, May 1952. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f50-023.

Hourston:1952:FGM

[32] Alan S. Hourston. The food and growth of the maskinonge (Esox masquinongy Mitchill) in Canadian Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(5):347–368, May 1952. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f50-024.

Quayle:1952:RDF

[33] D. B. Quayle. The relation of digging frequency to productivity of the butter clam, Saxidomus giganteus deshayes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(5):369–373, May 1952. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f50-025.

Manzer:1952:NDG

[34] J. I. Manzer. Notes on dispersion and growth of some British Columbia bottom fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(5): 374–377, May 1952. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f50-026.

Tully:1952:ODW

[35] John P. Tully. Oceanographic data of the Western Canadian Arctic region, 1935–37. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(5):

378–382, May 1952. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f50-027.

Smith:1952:LTA

[36] M. W. Smith. Limnology and trout angling in Charlotte County Lakes, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(6): 381–452, June 1952. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f50-028.

Brett:1952:SRS

[37] J. R. Brett. Skeena River sockeye escapement and distribution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(7):453–468, July 1952. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f50-029.

Keleher:1952:GTP

[38] J. J. Keleher. Growth and triaenophorus parasitism in relation to taxonomy of Lake Winnipeg ciscoes (Leucichthys). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(7):469–478, July 1952. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f50-030.

Manzer:1952:ETU

[39] J. I. Manzer. The effects of tagging upon a Pacific Coast flounder, Parophrys vetulus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(7): 479–485, July 1952. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f50-031.

Wilder:1952:RTC

[40] D. G. Wilder. The relative toxicity of certain metals to lobsters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(7):486–487, July 1952. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f50-032.

Berkeley:1952:RDP

[41] E. Berkeley and C. Berkeley. Re-discovery of the polychaete worm Trypanosyllis ingens Johnson. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 8c(7):488–490, July 1952. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f50-033.

Sprules:1952:ACW

[42] Wm. M. Sprules. The Arctic char of the West Coast of Hudson Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(1):1–15, January 1952. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f52-001.

Smith:1952:FOV

[43] Verna Z. Smith. Further ostracoda of the Vancouver Island region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(1):16–41, January 1952. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f52-002.

Doe:1952:CNT

[44] L. A. E. Doe. Currents and net transport in Loudoun Channel, April 1950. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(1):42–64, January 1952. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f52-003.

Dunbar:1952:SLC

[45] M. J. Dunbar and E. H. Grainger. Station list of the “Calanus” Expeditions, 1947–50. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(2): 65–82, February 1952. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f52-004.

Dunbar:1952:CSF

[46] M. J. Dunbar and H. H. Hildebrand. Contribution to the study of the fishes of Ungava Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(2):83– 128, February 1952. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f52-005.

Guttmann:1952:VBA

[47] A. Guttmann and F. A. Vandenheuvel. The vitamin B_{12} activity of natural material with particular reference to cod liver protein. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(3):129–140, March 1952. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f52-006.

Clemens:1952:FRS

[48] W. A. Clemens and G. V. Wilby. The first record of the smooth lumpsucker, Cyclopterichthys ventricosus (Pallas), in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(3):141–142, March 1952. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f52-007.

Barraclough:1952:AFP

[49] W. E. Barraclough. The agonid fish Pallasina barbata aix (Starks), from British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(3): 143–147, March 1952. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f52-008.

Castell:1952:IFF

[50] C. H. Castell and E. G. Mapplebeck. The importance of Flavobacterium in fish spoilage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(3):148– 156, March 1952. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f52-009.

Dussault:1952:IMR

[51] H. P. Dussault and R. A. Lachance. Improved medium for red halophilic bacteria from salt fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(3): 157–163, March 1952. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f52-010.

Maxwell:1952:DVB

[52] B. E. Maxwell. The distribution of vitamin B_{12}-active substances in some marine invertebrates of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(3):164–168, March 1952. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f52-011.

Wilder:1952:CSA

[53] D. G. Wilder. A comparative study of anadromous and freshwater populations of brook trout (Salvelinus fontinalis (Mitchill)). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(4):169–203, April 1952. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f52-012.

Ricker:1952:BCL

[54] W. E. Ricker. The benthos of Cultus Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(4):204–212, April 1952. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f52-013.

MacGregor:1952:CMG

[55] D. G. MacGregor and H. J. McLellan. Current measurements in the Grand Manan Channel. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(5):213– 222, May 1952. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f52-014.

Kielhorn:1952:BSZ

[56] William V. Kielhorn. The biology of the surface zone zooplankton of a boreo-Arctic Atlantic Ocean area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(5):223–264, May 1952. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f52-015.

Brett:1952:TTY

[57] J. R. Brett. Temperature tolerance in young Pacific salmon, genus Oncorhynchus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(6):265– 323, June 1952. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f52-016.

Hachey:1952:VTD

[58] H. B. Hachey. Vertical temperature distribution in the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(7):325–328, July 1952. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f52-017.

Millar:1952:STG

[59] F. Graham Millar. Surface temperatures of the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(7):329–394, July 1952. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f52-018.

Castell:1952:SRH

[60] C. H. Castell and E. G. Mapplebeck. The survival of red halophiles in water and in brines. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(8): 377–387, August 1952. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f52-019.

Fougere:1952:WTC

[61] H. Fougère. The water transfer in codfish muscle immersed in sodium chloride solutions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(8):388– 392, August 1952. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f52-020.

Khan:1952:SLF

[62] M. M. R. Khan. A study of the “Lipoxidase” in the flesh of British Columbia herring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9(8):393– 416, August 1952. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f52-021.

Carl:1953:LCL

[63] G. Clifford Carl. Limnobiology of Cowichan Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9a(9):417–449, September 1953. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f52-022.

Neave:1953:PAS

[64] Ferris Neave. Principles affecting the size of pink and chum salmon populations in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 9a(9):450–491, September 1953. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f52-023.

Huntsman:1953:MDL

[65] A. G. Huntsman. Movements and decline of large quoddy herring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(1):1–50, January 1953. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f53-001.

Kennedy:1953:MCF

[66] W. A. Kennedy. The morphometry of the Coregonine fishes of Great Bear Lake, N.W.T. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(2): 51–61, February 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f53-002.

Margolis:1953:EFB

[67] Leo Margolis. The effect of fasting on the bacterial flora of the intestine of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(2):62–63, February 1953. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f53-003.

Southcott:1953:VBC

[68] B. A. Southcott and H. L. A. Tarr. The vitamin B_{12} content of certain fishery materials. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(2): 64–68, February 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f53-004.

Tarr:1953:AHB

[69] H. L. A. Tarr. The action of hydroxylamine on bacteria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(2):69–75, February 1953. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f53-005.

S:1953:CNS

[70] Ira E. Cornwall F. G. S. The central nervous system of barnacles (Cirripedia). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(2):76–84, February 1953. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f53-006.

Berkeley:1953:MNS

[71] E. Berkeley and C. Berkeley. Micronereis nanaimoensis sp. n: with some notes on its life-history. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(2):

85–95, February 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f53-007.

Ketchum:1953:EFS

[72] Bostwick H. Ketchum and D. Jean Keen. The exchanges of fresh and salt waters in the Bay of Fundy and in Passamaquoddy Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(3):97–124, March 1953. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f53-008.

Pickard:1953:SVT

[73] G. L. Pickard and D. C. McLeod. Seasonal variation of temperature and salinity of surface waters of the British Columbia Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(3):125–145, March 1953. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f53-009.

Lauzier:1953:SLS

[74] Louis Lauzier. The St. Lawrence spring run-off and summer salinities in the Magdalen Shallows. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(3):146–147, March 1953. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f53-010.

Hachey:1953:WIS

[75] H. B. Hachey. A winter incursion of slope water on the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(3):148–153, March 1953. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f53-011.

McLellan:1953:SWS

[76] H. J. McLellan, L. Lauzier, and W. B. Bailey. The slope water off the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(4): 155–176, April 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f53-012.

Nutt:1953:CAO

[77] David C. Nutt. Certain aspects of oceanography in the coastal waters of Labrador. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(4):177– 186, April 1953. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f53-013.

Sullivan:1953:SFS

[78] Charlotte M. Sullivan and Kenneth C. Fisher. Seasonal fluctuations in the selected temperature of speckled trout, Salvelinus fontinalis (Mitchill). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(4):187–195, April 1953. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f53-014.

Black:1953:ULT

[79] Edgar C. Black. Upper lethal temperatures of some British Columbia freshwater fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(4): 196–210, April 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f53-015.

Riley:1953:TGC

[80] Gordon A. Riley. Theory of growth and competition in natural populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(5):211– 223, May 1953. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f53-016.

Rawson:1953:SCN

[81] D. S. Rawson. The standing crop of net plankton in lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(5):224–237, May 1953. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f53-017.

Langford:1953:MPC

[82] R. R. Langford. Methods of plankton collection and a description of a new sampler. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(5): 238–252, May 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f53-018.

McCombie:1953:FIG

[83] A. M. McCombie. Factors influencing the growth of phytoplankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(5):253–282, May 1953. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f53-019.

Kleerekoper:1953:MP

[84] Herman Kleerekoper. The mineralization of plankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(5):283–291, May 1953. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f53-020.

Foerster:1953:CSC

[85] R. E. Foerster and W. E. Ricker. The Coho salmon of Cultus Lake and Sweltzer Creek. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(6): 293–319, June 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f53-021.

Pike:1953:CPH

[86] Gordon C. Pike. Colour pattern of humpback whales from the coast of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(6): 320–325, June 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f53-022.

Grainger:1953:AGM

[87] E. H. Grainger. On the age, growth, migration, reproductive potential and feeding habits of the Arctic char (Salvelinus alpinus) of Frobisher Bay, Baffin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(6): 326–370, June 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f53-023.

Wilder:1953:GRA

[88] D. G. Wilder. The growth rate of the American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(7):371– 412, July 1953. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f53-024.

Kennedy:1953:GMF

[89] W. A. Kennedy. Growth, maturity, fecundity and mortality in the relatively unexploited whitefish, Coregonus clupeaformis, of Great Slave Lake.

Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(7):413–441, July 1953. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f53-025.

Gordon:1953:EAO

[90] H. Scott Gordon. An economic approach to the optimum utilization of fishery resources. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(7): 442–457, July 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f53-026.

Ketchen:1953:UCE

[91] K. S. Ketchen. The use of catch–effort and tagging data in estimating a flatfish population. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(8): 459–485, August 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f53-027.

Rawson:1953:BFG

[92] D. S. Rawson. The bottom fauna of Great Slave Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(8):486–520, August 1953. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f53-028.

Lusena:1953:PVC

[93] C. V. Lusena and R. C. Rose. Preparation and viscosity of chitosans. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(8):521–522, August 1953. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f53-029.

MacKinnon:1953:RCC

[94] Dixon MacKinnon and William S. Hoar. Responses of Coho and chum salmon fry to current. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(8): 523–538, August 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f53-030.

Scott:1953:EHS

[95] D. M. Scott. Experiments with the harbour seal, Phoca vitulina, a definitive host of a marine nematode, Porrocaecum decipiens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(8):539–547, August 1953. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f53-031.

Brett:1953:PEU

[96] J. R. Brett and D. MacKinnon. Preliminary experiments using lights and bubbles to deflect migrating young spring salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(8):548–559, August 1953. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f53-032.

Mead:1953:COF

[97] Giles W. Mead and F. H. C. Taylor. A collection of oceanic fishes from off Northeastern Japan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(8): 560–582, August 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f53-033.

Zimmerman:1953:USB

[98] Howard K. Zimmerman, Jr. The ultraviolet spectrum and basicity of “acid-soluble” chitin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(8): 583–589, August 1953. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f53-034.

Stern:1953:SPE

[99] Joseph A. Stern and Ernest E. Lockhart. A study of the proteolytic enzyme activity of the pyloric caeca of redfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 10(8):590–598, August 1953. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f53-035.

Lawler:1954:OTP

[100] G. H. Lawler. Observations on the trout-perch Percopsis omiscomaycus (Walbaum), at Heming Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(1):1–4, January 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-001.

Martin:1954:CWF

[101] N. V. Martin. Catch and winter food of lake trout in certain Algonquin Park lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(1):5–10, January 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f54-002.

Templeman:1954:ELF

[102] Wilfred Templeman and A. M. Fleming. European ling — first record from North American waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(1):11–13, January 1954. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f54-003.

Herlinveaux:1954:STC

[103] R. H. Herlinveaux. Surface tidal currents in Juan de Fuca Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(1):14–31, January 1954. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f54-004.

Bailey:1954:AVT

[104] W. B. Bailey, D. G. MacGregor, and H. B. Hachey. Annual variations of temperature and salinity in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(1):32–47, January 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-005.

Mackay:1954:TCO

[105] Bruce S. Mackay. Tidal current observations in Hecate Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(1):48–56, January 1954. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f54-006.

Jampolsky:1954:GHS

[106] A. Jampolsky and W. S. Hoar. Growth hormone from salmon pituitary glands. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(1):57–62, January 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f54-007.

Potter:1954:PAC

[107] Gilbert D. Potter and William S. Hoar. The presence of androgens in chum salmon (Oncorhynchus keta Walbaum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(1):63–68, January 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-008.

Hoar:1954:BJP

[108] William S. Hoar. The behaviour of juvenile Pacific salmon, with particular reference to the sockeye (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(1):69–97, January 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-009.

Grainger:1954:SLC

[109] E. H. Grainger. Station list of the Calanus Expeditions, 1951–52, together with Frobisher Bay stations, 1948, 1950 and 1951, and Resolution Island stations, 1950. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(1):98– 105, January 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f54-010.

Hasler:1954:OPO

[110] Arthur D. Hasler. Odour perception and orientation in fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(2):107–129, February 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f54-011.

Kleerekoper:1954:HFS

[111] H. Kleerekoper and E. C. Chagnon. Hearing in fish, with special reference to Semotilus atromaculatus atromaculatus (Mitchill). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(2):130–152, February 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-012.

Sullivan:1954:TRR

[112] Charlotte M. Sullivan. Temperature reception and responses in fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(2):153–170, February 1954.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f54-013.

Scott:1954:CSY

[113] D. M. Scott. A comparative study of the yellowtail flounder from three Atlantic fishing areas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(3): 171–197, March 1954. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f54-014.

Huntsman:1954:GOS

[114] A. G. Huntsman, W. B. Bailey, and H. B. Hachey. The general oceanography of the Strait of Belle Isle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(3):198–260, March 1954. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f54-015.

Dussault:1954:ADP

[115] H. P. Dussault. Adaptation of the “drop plate” method for the enumeration of red halophilic bacteria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(3):261–266, March 1954. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f54-016.

Margolis:1954:LPR

[116] L. Margolis. List of the parasites recorded from sea mammals caught off the West Coast of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(3):267–283, March 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-017.

Kennedy:1954:TRA

[117] W. A. Kennedy. Tagging returns, age studies and fluctuations in abundance of Lake Winnipeg whitefish, 1931–1951. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(3):284–309, March 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-018.

Brett:1954:SAO

[118] J. R. Brett and D. MacKinnon. Some aspects of olfactory perception in migrating adult Coho and spring salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(3):310–318, March 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-019.

Margolis:1954:TKW

[119] L. Margolis. Three kinds of whale-lice (Cyamidae: Amphipoda) from the Pacific Coast of Canada, including a new species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(3):319–325, March 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-020.

Berkeley:1954:NLH

[120] E. Berkeley and C. Berkeley. Notes on the life-history of the polychaete, Dodecaceria fewkesi (nom. n.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(3):326–334, March 1954. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f54-021.

Ketchen:1954:RSR

[121] K. S. Ketchen. The rockfish Sebastodes rubrivinctus in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(3):335–338, March 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f54-022.

Foerster:1954:RAS

[122] R. E. Foerster. On the relation of adult sockeye salmon (Oncorhynchus nerka) returns to known smolt seaward migrations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(4):339–350, April 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-023.

Templeman:1954:MSD

[123] Wilfred Templeman. Migrations of spiny dogfish tagged in Newfoundland waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(4):351– 354, April 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f54-024.

Truscott:1954:PCL

[124] Beryl Truscott, D. G. Gage, and P. L. Hoogland. Preparation of cod liver residues and vitamin B_{12} concentrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(4):355–361, April 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-025.

McCrimmon:1954:SSP

[125] H. R. McCrimmon. Stream studies on planted Atlantic salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(4):362–403, April 1954. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f54-026.

McLellan:1954:BTS

[126] H. J. McLellan. Bottom temperatures on the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(4):404–418, April 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-027.

McLellan:1954:TSR

[127] H. J. McLellan. Temperature-salinity relations and mixing on the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(4):419– 430, April 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f54-028.

Udvardy:1954:DAR

[128] Miklos D. F. Udvardy. Distribution of appendicularians in relation to the Strait of Belle Isle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(4): 431–453, April 1954. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f54-029.

Berkeley:1954:APF

[129] E. Berkeley and C. Berkeley. Additions to the polychaete fauna of Canada, with comments on some older records. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(4):454–471, April 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-030.

Wisby:1954:EOO

[130] Warren J. Wisby and Arthur D. Hasler. Effect of olfactory occlusion on migrating silver salmon (O. kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(4):472–478, April 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-031.

McMahon:1954:ADE

[131] V. H. McMahon. The abundance and distribution of entomostracan plankton at Lakelse Lake, B.C., 1949–1952. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(4):479–499, April 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-032.

Barraclough:1954:OCD

[132] W. E. Barraclough and M. Waldichuk. Offshore characteristics in the deep waters of the Strait of Georgia as indicated by bathypelagic fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(5):501–506, May 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f54-033.

Grainger:1954:PAU

[133] E. H. Grainger. Polychaetous annelids of Ungava Bay, Hudson Strait, Frobisher Bay and Cumberland Sound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(5):507–528, May 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-034.

Harrington:1954:CST

[134] Robert W. Harrington, Jr. Contrasting susceptibilities of two fish species to a diet destructive to vitamin B_{1}. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(5):529–534, May 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-035.

Homans:1954:RBF

[135] R. E. S. Homans and V. D. Vladykov. Relation between feeding and the sexual cycle of the haddock. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(5):535–542, May 1954. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f54-036.

Jewett:1954:NSP

[136] Stanley G. Jewett, Jr. New stoneflies (Plecoptera) from Western North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(5):543– 549, May 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f54-037.

Miller:1954:MCT

[137] Richard B. Miller. Movements of cutthroat trout after different periods of retention upstream and downstream from their homes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(5):550–558, May 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-038.

Ricker:1954:SR

[138] W. E. Ricker. Stock and recruitment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(5):559–623, May 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-039.

Robertson:1954:TSP

[139] J. G. Robertson. The trophic status of Port John Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(5):624–651, May 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f54-040.

Slastenenko:1954:RGH

[140] E. P. Slastenenko. The relative growth of hybrid char (Salvelinus fontinalis × Cristivomer namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(5):652–659, May 1954. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f54-041.

Stevenson:1954:AGR

[141] J. A. Stevenson and L. M. Dickie. Annual growth rings and rate of growth of the giant scallop, Placopecten magellanicus (Gmelin) in the Digby area of the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(5): 660–671, May 1954. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f54-042.

Bangham:1954:SPF

[142] Ralph V. Bangham and James R. Adams. A survey of the parasites of freshwater fishes from the Mainland of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):673–708, June 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-043.

Dunbar:1954:ACU

[143] M. J. Dunbar. The amphipod crustacea of Ungava Bay, Canadian Eastern Arctic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):709– 798, June 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f54-044.

Herlinveaux:1954:TCJ

[144] R. H. Herlinveaux. Tidal currents in Juan de Fuca Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):799–815, June 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-045.

Humes:1954:TCS

[145] Arthur G. Humes. Tisbe celata n. sp., a harpacticoid copepod from the mantle cavity of the edible mussel in New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):816–826, June 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-046.

Kennedy:1954:GMM

[146] W. A. Kennedy. Growth, maturity and mortality in the relatively unexploited lake trout, Cristivomer namaycush, of Great Slave Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):827–852, June 1954. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f54-047.

LaCroix:1954:AMS

[147] G. Wilfred LaCroix and John P. Tully. The anomaly of mean sea level in Seymour narrows, B.C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):853–883, June 1954. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f54-048.

Lawler:1954:NGD

[148] G. H. Lawler and W. B. Scott. Notes on the geographical distribution and the hosts of the cestode genus Triaenophorus in North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):884–893, June 1954. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f54-049.

Scott:1954:EIA

[149] D. M. Scott. Experimental infection of Atlantic cod with a larval marine nematode from smelt. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6): 894–900, June 1954. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f54-050.

Vaisey:1954:OSE

[150] E. B. Vaisey. Osmophilism of Sporendonema epizoum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):901–903, June 1954. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f54-051.

Vladykov:1954:TCE

[151] Vadim D. Vladykov. Taxonomic characters of the Eastern North America chars (Salvelinus and Cristivomer). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):904–932, June 1954. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f54-052.

Wickett:1954:OSS

[152] W. Percy Wickett. The oxygen supply to salmon eggs in spawning beds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):933–953, June 1954. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f54-053.

Wolfgang:1954:STSa

[153] R. W. Wolfgang. Studies of the trematode Stephanostomum baccatum (Nicoll, 1907): I. The distribution of the metacercaria in Eastern Canadian flounders. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6): 954–962, June 1954. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f54-054.

Wolfgang:1954:STSb

[154] R. W. Wolfgang. Studies of the trematode Stephanostomum baccatum (Nicoll, 1907): II. Biology, with special reference to the stages affecting the winter flounder. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6): 963–987, June 1954. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f54-055.

Foerster:1954:SRS

[155] R. E. Foerster. Sex ratios in sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 11(6):988–997, June 1954. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f54-056.

Doe:1955:OWC

[156] L. A. E. Doe. Offshore waters of the Canadian Pacific Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):1–34, January 1955. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f55-001.

Partlo:1955:DAG

[157] J. M. Partlo. Distribution, age and growth of Eastern Pacific albacore (Thunnus alalunga Gmelin). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):35–60, January 1955. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f55-002.

Partlo:1955:HSA

[158] J. M. Partlo. Histological studies on albacore (Thunnus alalunga) gonads from the Eastern Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):61–67, January 1955. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f55-003.

Legendre:1955:ADL

[159] R. Legendre. The artificial drying of lightly salted codfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):68–74, January 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-004.

Dingle:1955:ESP

[160] J. R. Dingle, D. E. Eagles, and J. M. Neelin. Electrophoretic studies of proteins extractable from post-rigor cod at low ionic strength. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):75–84, January 1955. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f55-005.

Gordon:1955:EFC

[161] H. Scott Gordon. Economic factors in catch fluctuations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):85–92, January 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-006.

Dombroski:1955:CNI

[162] E. Dombroski. Cestode and nematode infection of sockeye smolts from Babine Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):93–96, January 1955. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f55-007.

Margolis:1955:SHP

[163] Leo Margolis and Gordon C. Pike. Some helminth parasites of Canadian Pacific whales. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):97– 120, January 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f55-008.

Margolis:1955:NMT

[164] Leo Margolis. Notes on the morphology, taxonomy and synonymy of several species of whale-lice (Cyamidae: Amphipoda). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):121–133, January 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-009.

Forrester:1955:RSW

[165] C. R. Forrester and K. S. Ketchen. The resistance to salt water corrosion of various types of metal wire used in the tagging of flatfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):134–142, January 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-010.

Andrews:1955:MFC

[166] E. A. Andrews. More folliculinids (Ciliata heterotricha) from British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):143–146, January 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f55-011.

Fraser:1955:SBF

[167] J. M. Fraser. The smallmouth bass fishery of South Bay, Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):147–177, January 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f55-012.

Hoar:1955:PPR

[168] William S. Hoar. Phototactic and pigmentary responses of sockeye salmon smolts following injury to the pineal organ. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(1):178–185, January 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-013.

Dickie:1955:IDP

[169] L. M. Dickie and F. D. McCracken. Isopleth diagrams to predict equilibrium yields of a small flounder fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(2):187–209, February 1955. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f55-014.

Smith:1955:FPC

[170] M. W. Smith. Fertilization and predator control to improve trout angling in natural lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(2): 210–237, February 1955. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f55-015.

Smith:1955:AEC

[171] M. W. Smith and J. W. Saunders. The American eel in certain fresh waters of the Maritime Provinces of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(2):238–269, February 1955. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f55-016.

Medcof:1955:DNC

[172] J. C. Medcof. Day and night characteristics of spatfall and of behaviour of oyster larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(2): 270–286, February 1955. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f55-017.

Giglioli:1955:EMN

[173] M. E. C. Giglioli. The egg masses of the Naticidae (Gastropoda). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(2):287–327, February 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f55-018.

Castell:1955:SHT

[174] C. H. Castell and Rosalie E. Triggs. Spoilage of haddock in the trawlers at sea: The measurement of spoilage and standards of quality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(3):329–341, March 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-019.

Bousfield:1955:SPF

[175] E. L. Bousfield. Some physical features of the Miramichi Estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(3):342–361, March 1955. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f55-020.

MacKinnon:1955:SOM

[176] D. MacKinnon and J. R. Brett. Some observations on the movement of Pacific salmon fry through a small impounded water basin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(3):362–368, March 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-021.

Neave:1955:NSM

[177] Ferris Neave. Notes on the seaward migration of pink and chum salmon fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(3):369–374, March 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f55-022.

McLellan:1955:CBT

[178] H. J. McLellan. Change in bottom temperatures on the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(3):375–386, March 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f55-023.

Shepard:1955:RTY

[179] M. P. Shepard. Resistance and tolerance of young speckled trout (Salvelinus fontinalis) to oxygen lack, with special reference to low oxygen acclimation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(3):387– 446, March 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f55-024.

Parker:1955:MRE

[180] Richard A. Parker. A method for removing the effect of recruitment on Petersen-type population estimates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(3):447–450, March 1955. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f55-025.

Milne:1955:SRS

[181] D. J. Milne. The Skeena River salmon fishery, with special reference to sockeye salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(3): 451–485, March 1955. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f55-026.

Martin:1955:LBO

[182] N. V. Martin. Limnological and biological observations in the region of the Ungava or Chubb Crater, Province of Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(4):487–498, April 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-027.

Godfrey:1955:ESR

[183] H. Godfrey. On the ecology of Skeena River whitefishes, Coregonus and Prosopium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(4):499– 542, April 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f55-028.

Carlander:1955:SCF

[184] Kenneth D. Carlander. The standing crop of fish in lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(4):543–570, April 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-029.

Bishop:1955:SEA

[185] Yvonne M. M. Bishop and Leo Margolis. A statistical examination of Anisakis larvae (Nematoda) in herring (Clupea pallasi) of the British Columbia Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(4): 571–592, April 1955. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f55-030.

Lea:1955:CWC

[186] Helen E. Lea. The chaetognaths of Western Canadian coastal waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(4):593–617, April 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f55-031.

Hayes:1955:TPN

[187] F. R. Hayes and D. A. Livingstone. The trout population of a Nova Scotia lake as affected by habitable water, poisoning of the shallows and stocking. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(4):618–635, April 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f55-032.

Guttmann:1955:EEH

[188] A. Guttmann. Extracts and enzymic hydrolysates from fish liver and mammalian liver. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5): 637–645, May 1955. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f55-033.

Dingle:1955:NAP

[189] J. R. Dingle and W. J. Dyer. Note on albumin protein fractions in a sturgeon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5):646– 648, May 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f55-034.

Tester:1955:ERN

[190] Albert L. Tester. Estimation of recruitment and natural mortality rate from age-composition and catch data in British Columbia herring populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5):649– 681, May 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f55-035.

Vavasour:1955:PSH

[191] G. R. Vavasour and A. A. Blair. Pollution survey of Humber Arm and Exploits River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5): 682–697, May 1955. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f55-036.

McKenzie:1955:ACP

[192] R. A. McKenzie and G. F. M. Smith. Atlantic cod populations along the Southern Canadian Mainland as shown by vertebral count studies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5):698–705, May 1955. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f55-037.

Pollard:1955:MST

[193] R. A. Pollard. Measuring seepage through salmon spawning gravel. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5):706–741, May 1955. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f55-038.

Smith:1955:RBS

[194] S. B. Smith. The relation between scale diameter and body length of Kamloops trout, Salmo gairdneri kamloops. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5):742–753, May 1955. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f55-039.

Vladykov:1955:CAS

[195] Vadim D. Vladykov. A comparison of Atlantic sea sturgeon with a new subspecies from the Gulf of Mexico (Acipenser oxyrhynchus de sotoi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5):754–761, May 1955.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f55-040.

Bousfield:1955:CCH

[196] E. L. Bousfield. The cirripede crustacea of the Hudson Strait region, Canadian Eastern Arctic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5):762–767, May 1955. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f55-041.

Fuller:1955:ISL

[197] W. A. Fuller. The inconnu (Stenodus leucichthys mackenziei) in Great Slave Lake and adjoining waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5):768–780, May 1955. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f55-042.

Tuck:1955:FFH

[198] Leslie M. Tuck and H. J. Squires. Food and feeding habits of Brünnich’s murre (Uria lomvia lomvia) on akpatok Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(5):781–792, May 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-043.

Anonymous:1955:O

[199] Anonymous. Obituary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(6): 794–796, June 1955. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f55-044.

Dickie:1955:FAG

[200] L. M. Dickie. Fluctuations in abundance of the giant scallop, Placopecten magellanicus (Gmelin), in the Digby area of the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(6):797–857, June 1955. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f55-045.

Fontaine:1955:PCC

[201] Marion Fontaine. The planktonic copepods (Calanoida, Cyclopoida, Monstrilloida) of Ungava Bay, with special reference to the biology of Pseudocalanus minutus and Calanus finmarchicus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(6):858–898, June 1955. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f55-046.

Grainger:1955:EUB

[202] E. H. Grainger. Echinoderms of Ungava Bay, Hudson Strait, Frobisher Bay and Cumberland Sound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(6):899–916, June 1955. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f55-047.

Black:1955:BLH

[203] Edgar C. Black. Blood levels of hemoglobin and lactic acid in some freshwater fishes following exercise. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(6):917–929, June 1955. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f55-048.

Hewson:1955:AMS

[204] L. C. Hewson. Age, maturity, spawning and food of burbot, Lota lota, in Lake Winnipeg. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(6): 930–940, June 1955. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f55-049.

Anonymous:1955:CRP

[205] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(6): 941–942, June 1955. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f55-050.

Anonymous:1955:IVX

[206] Anonymous. Index to Volume XII of the Journal of the Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 12(6): 943–950, June 1955. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f55-051.

Fry:1956:MDC

[207] F. E. J. Fry. Movements of drift cards in Georgian Bay in 1953. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(1):1–5, January 1956. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f56-001.

Waldichuk:1956:BPT

[208] Michael Waldichuk. Basic productivity of Trevor Channel and Alberni Inlet from chemical measurements. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(1):7–20, January 1956. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f56-002.

Castell:1956:SFVa

[209] C. H. Castell, W. A. MacCallum, and H. E. Power. Spoilage of fish in the vessels at sea: 2. Treatment on the deck and in the hold. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(1):21–39, January 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-003.

Grainger:1956:SLC

[210] E. H. Grainger and M. J. Dunbar. Station list of the “Calanus” Expeditions, 1953–4. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(1): 41–45, January 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f56-004.

Morton:1956:FO

[211] Margaret L. Morton and W. J. Dyer. Frozen oysters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(1):47–51, January 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-005.

Rawson:1956:NPG

[212] D. S. Rawson. The net plankton of Great Slave Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(1):53–127, January 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-006.

Dyer:1956:SFP

[213] W. J. Dyer and Margaret L. Morton. Storage of frozen plaice fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(1):129–134, January 1956. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f56-007.

Miller:1956:CRS

[214] Richard B. Miller. The collapse and recovery of a small whitefish fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(1):135–146, January 1956. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f56-008.

Templeman:1956:JCA

[215] Wilfred Templeman and Gertrude L. Andrews. Jellied condition in the American plaice Hippoglossoides platessoides (Fabricius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(2):147–182, February 1956. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f56-009.

Dussault:1956:SRHa

[216] H. P. Dussault. Study of red halophilic bacteria in solar salt and salted fish: I. Effect of bacto-oxgall. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(2):183–194, February 1956. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f56-010.

Dussault:1956:SRHb

[217] H. P. Dussault. Study of red halophilic bacteria in solar salt and salted fish: II. Bacto-oxgall as a selective agent for differentiation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(2):195–199, February 1956. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f56-011.

Castell:1956:NPN

[218] C. H. Castell and E. G. Mapplebeck. A note on the production of nitrite from hydroxylamine by some heterotrophic bacteria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(2):201–206, February 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-012.

Castell:1956:SFVb

[219] C. H. Castell and G. K. Gunnarsson. Spoilage of fish in the vessels at sea: 3. The value of nitrite ice and nitrite dips for the preservation of gutted fish in the hold of the vessel. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(2):207–218, February 1956. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f56-013.

Blair:1956:AST

[220] A. A. Blair. Atlantic salmon tagged in East Coast Newfoundland waters at Bonavista. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(2):219–232, February 1956. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f56-014.

Berkeley:1956:CPA

[221] E. Berkeley and C. Berkeley. On a collection of polychaetous annelids from Northern Banks Island, from the South Beaufort Sea, and from Northwest Alaska; together with some new records from the East Coast of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(2):233–246, February 1956. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f56-015.

McLeese:1956:ETS

[222] D. W. McLeese. Effects of temperature, salinity and oxygen on the survival of the American lobster. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(2):247–272, February 1956. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f56-016.

Kirsch:1956:IRS

[223] Milton Kirsch. Ionic ratios of some of the major components in riverdiluted sea water in Bute and Knight Inlets, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3):273–289, March 1956. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f56-017.

Castell:1956:SFVc

[224] C. H. Castell and Maxine F. Greenough. Spoilage of fish in the vessels at sea: 4. Effect of removal of gills on rate of spoilage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3):291–296, March 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-018.

McLellan:1956:SOB

[225] H. J. McLellan. On the sharpness of oceanographic boundaries south of Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3): 297–301, March 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f56-019.

Bailey:1956:ODB

[226] W. B. Bailey. On the origin of deep Baffin Bay water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3):303–308, March 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-020.

Hoar:1956:BMP

[227] William S. Hoar. The behaviour of migrating pink and chum salmon fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3):309–325, March 1956. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f56-021.

Larkin:1956:ICP

[228] P. A. Larkin. Interspecific competition and population control in freshwater fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3):327– 342, March 1956. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f56-022.

Scott:1956:SIL

[229] D. M. Scott. On the specific identity of the larval Porrocaecum (Nematoda) in Atlantic cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3): 343–356, March 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f56-023.

Ketchen:1956:CTF

[230] K. S. Ketchen. Climatic trends and fluctuations in yield of marine fisheries of the Northeast Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3):357–374, March 1956. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f56-024.

Kleerekoper:1956:IES

[231] H. Kleerekoper and Kira Sibakin. An investigation of the electrical “spike” potentials produced by the sea lamprey (Petromyzon marinus) in the water surrounding the head region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3):375–383, March 1956. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f56-025.

Trites:1956:OCS

[232] Ronald W. Trites. The oceanography of Chatham Sound, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3):385– 434, March 1956. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f56-026.

MacGregor:1956:CTC

[233] D. G. MacGregor. Currents and transport in Cabot Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(3):435–448, March 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-027.

Tibbo:1956:PHC

[234] S. N. Tibbo. Populations of herring (Clupea harengus L.) in Newfoundland waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4):449– 466, April 1956. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f56-028.

Templeman:1956:ROL

[235] Wilfred Templeman and H. J. Squires. Relationship of otolith lengths and weights in the haddock Melanogrammus aeglefinus (L.) to the rate of growth of the fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4): 467–487, April 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f56-029.

Margolis:1956:PHA

[236] Leo Margolis. Parasitic helminths and arthropods from pinnipedia of the Canadian Pacific Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4):489–505, April 1956. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f56-030.

Fabris:1956:DID

[237] David A. Fabris. Diet-induced differences in the weight — length relationship of aquarium fed sardines (Sardinops caerulea). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4):507–513, April 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-031.

Northcote:1956:IPB

[238] T. G. Northcote and P. A. Larkin. Indices of productivity in British Columbia lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4): 515–540, April 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f56-032.

Berkeley:1956:NPE

[239] E. Berkeley and C. Berkeley. Notes on Polychaeta from the East Coast of Vancouver Island and from adjacent waters, with a description of a new species of Aricidea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4): 541–546, April 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f56-033.

Lapworth:1956:EFP

[240] E. D. Lapworth. The effect of fry plantings on whitefish production in Eastern Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4): 547–558, April 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f56-034.

Vaisey:1956:CCN

[241] E. B. Vaisey. Chemical changes in nitrite-treated Atlantic cod fillets in relation to spoilage assessment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4):559–567, April 1956. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f56-035.

Dyer:1956:SFR

[242] W. J. Dyer, Margaret L. Morton, Doris I. Fraser, and E. G. Bligh. Storage of frozen rosefish fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4): 569–579, April 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f56-036.

Pickard:1956:SBC

[243] G. L. Pickard. Surface and bottom currents in the Strait of Georgia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4):581–590, April 1956. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f56-037.

Ellis:1956:STS

[244] D. V. Ellis. Some temperature and salinity records from the Canadian Arctic during 1954 and 1955. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(4):591–598, April 1956. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f56-038.

Crossman:1956:GMM

[245] E. J. Crossman. Growth, mortality and movements of a sanctuary population of maskinonge (Esox masquinongy Mitchill). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(5):599–612, May 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-039.

Watt:1956:CSM

[246] Kenneth E. F. Watt. The choice and solution of mathematical models for predicting and maximizing the yield of a fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(5):613–645, May 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-040.

Ketchen:1956:FIS

[247] K. S. Ketchen. Factors influencing the survival of the lemon sole (Parophrys vetulus) in Hecate Strait, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(5):647–693, May 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-041.

Johnson:1956:DYS

[248] W. E. Johnson. On the distribution of young sockeye salmon (Oncorhynchus nerka) in Babine and Nilkitkwa Lakes, B.C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(5):695–708, May 1956. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f56-042.

Nutt:1956:OHF

[249] David C. Nutt and Lawrence K. Coachman. The oceanography of Hebron Fjord, Labrador. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(5): 709–758, May 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f56-043.

Lindsey:1956:DTF

[250] C. C. Lindsey. Distribution and taxonomy of fishes in the Mackenzie Drainage of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(6):759–789, June 1956. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f56-044.

Dambergs:1956:AWM

[251] N. Dambergs. Acetone–water mixtures for the extraction and rapid estimation of fats of biological materials, particularly fish products. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(6):791–797, June 1956. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f56-045.

Vladykov:1956:FWS

[252] Vadim D. Vladykov. Fecundity of wild speckled trout (Salvelinus fontinalis) in Quebec lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(6): 799–841, June 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f56-046.

Andrews:1956:BAC

[253] C. W. Andrews and E. Lear. The biology of Arctic char (Salvelinus alpinus L.) in Northern Labrador. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(6):843–860, June 1956. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f56-047.

McLean:1956:SFV

[254] Norma L. McLean and C. H. Castell. Spoilage of fish in the vessels at sea: 5. Bilgy fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(6): 861–868, June 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f56-048.

Freeman:1956:PCH

[255] H. C. Freeman and P. L. Hoogland. Processing of cod and haddock viscera: 1. Laboratory experiments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(6):869–877, June 1956. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f56-049.

Anonymous:1956:IVJ

[256] Anonymous. Index to volume 13 of the Journal of the Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 13(6): 879–886, June 1956. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f56-050.

Hayes:1957:VBF

[257] F. R. Hayes. On the variation in bottom fauna and fish yield in relation to trophic level and lake dimensions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(1):1–32, January 1957. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f57-001.

Vallentyne:1957:MNO

[258] J. R. Vallentyne. The molecular nature of organic matter in lakes and oceans, with lesser reference to sewage and terrestrial soils. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(1):33–82, January 1957. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f57-002.

Sergeant:1957:SCE

[259] D. E. Sergeant and H. D. Fisher. The smaller cetacea of Eastern Canadian Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(1):83–115, January 1957. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f57-003.

Black:1957:ABLa

[260] Edgar C. Black. Alterations in the blood level of lactic acid in certain salmonoid fishes following muscular activity: I. Kamloops trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(2):117–134, February 1957. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f57-004.

Blair:1957:STF

[261] A. A. Blair. Salmon tagging at Francis Harbour Bight, Labrador. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(2):135–140, February 1957. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f57-005.

Blair:1957:STC

[262] A. A. Blair. Salmon tagging at Cape Charles, Labrador. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(2):141–144, February 1957. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f57-006.

Kleerekoper:1957:IES

[263] H. Kleerekoper and Kira Sibakin. An investigation of the electrical “spike” potentials produced by the sea lamprey (Petromyzon marinus) in the water surrounding the head region. II. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(2):145–151, February 1957. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f57-007.

Godfrey:1957:MAD

[264] Harold Godfrey. Mortalities among developing trout and salmon ova following shock by direct-current electrical fishing gear. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(2):153–164, February 1957. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f57-008.

Cameron:1957:SFI

[265] Frances Ella Cameron. Some factors influencing the distribution of pelagic copepods in the Queen Charlotte Islands area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(2):165–202, February 1957. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f57-009.

Vernon:1957:SCY

[266] E. H. Vernon and R. G. McMynn. Scale characteristics of yearling coastal cutthroat and steelhead trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(2):203–212, February 1957. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f57-010.

McLellan:1957:DOS

[267] H. J. McLellan. On the distinctness and origin of the slope water off the Scotian Shelf and its easterly flow south of the Grand Banks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(2):213–239, February 1957. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f57-011.

Tully:1957:PWS

[268] John P. Tully and Allan J. Dodimead. Properties of the water in the Strait of Georgia, British Columbia, and influencing factors. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(3):241–319, March 1957. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f57-012.

Waldichuk:1957:POS

[269] Michael Waldichuk. Physical oceanography of the Strait of Georgia, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(3): 321–486, March 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f57-013.

Tabata:1957:POB

[270] Susumu Tabata and George L. Pickard. The physical oceanography of Bute Inlet, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4):487–520, April 1957. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f57-014.

Legare:1957:QQD

[271] Joseph Eugène Henri Légaré. The qualitative and quantitative distribution of plankton in the Strait of Georgia in relation to certain oceanographic factors. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4): 521–552, April 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f57-015.

Roussow:1957:SCC

[272] George Roussow. Some considerations concerning sturgeon spawning periodicity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4):553– 572, April 1957. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f57-016.

Vernon:1957:MCT

[273] E. H. Vernon. Morphometric comparison of three races of Kokanee (Oncorhynchus nerka) within a large British Columbia lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4):573–598, April 1957. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f57-017.

Smith:1957:EDC

[274] Stanford H. Smith. Evolution and distribution of the coregonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4):599–604, April 1957. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f57-018.

Pickard:1957:FWT

[275] G. L. Pickard and R. W. Trites. Fresh water transport determination from the heat budget with applications to British Columbia inlets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4):605–616, April 1957. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f57-019.

Castell:1957:APFa

[276] C. H. Castell and Maxine F. Greenough. The action of Pseudomonas on fish muscle: 1. Organisms responsible for odours produced during incipient spoilage of chilled fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4):617–625, April 1957. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f57-020.

Dyer:1957:IIS

[277] W. J. Dyer, Doris I. Fraser, D. G. Ellis, and W. A. MacCallum. Influence of intermittent short storage periods at 15^{o}f., as encountered during refrigerator car transportation, on the quality of frozen cod stored at 0^{o}F. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4):625–635, April 1957. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f57-021.

Bligh:1957:YDD

[278] E. G. Bligh, W. J. Dyer, and D. C. Horne. Yellow discoloration and deterioration in frozen lobster meat. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4):637–644, April 1957. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f57-022.

Black:1957:ABLb

[279] Edgar C. Black. Alterations in the blood level of lactic acid in certain salmonoid fishes following muscular activity: II. Lake trout, Salvelinus namaycush. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4):645– 649, April 1957. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f57-023.

Lindsey:1957:PEW

[280] C. C. Lindsey. Possible effects of water diversions on fish distribution in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(4): 651–668, April 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f57-024.

Schaefer:1957:SCP

[281] Milner B. Schaefer. Some considerations of population dynamics and economics in relation to the management of the commercial marine fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(5):669–681, May 1957. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f57-025.

Frick:1957:OLF

[282] Harold C. Frick. The optimum level of fisheries exploitation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(5):683–686, May 1957. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f57-026.

Miller:1957:PSH

[283] Richard B. Miller. Permanence and size of home territory in streamdwelling cutthroat trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(5):687–691, May 1957. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f57-027.

Squires:1957:SII

[284] H. J. Squires. Squid, Illex illecebrosus (LeSueur), in the Newfoundland fishing area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(5):693– 728, May 1957. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f57-028.

Mills:1957:HGC

[285] D. H. Mills. Herring gulls and common terns as possible predators of lobster larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(5): 729–730, May 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f57-029.

Bailey:1957:OFC

[286] W. B. Bailey. Oceanographic features of the Canadian Archipelago. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(5):731–769, May 1957. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f57-030.

Castell:1957:ECA

[287] C. H. Castell and Maxine F. Greenough. The effect of certain antibiotics on the production of trimethylamine and hydrogen sulphide by bacterial enzymes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(5): 771–774, May 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f57-031.

Castell:1957:APFb

[288] C. H. Castell, Maxine F. Greenough, and Norma L. Jenkin. The action of Pseudomonas on fish muscle: 2. Musty and potato-like odours. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(5):775–782, May 1957. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f57-032.

Alderdice:1957:SEK

[289] D. F. Alderdice and J. R. Brett. Some effects of kraft mill effluent on young Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(5): 783–795, May 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f57-033.

Miller:1957:GPF

[290] Richard B. Miller. Have the genetic patterns of fishes been altered by introductions or by selective fishing? Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6):797–806, June 1957. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f57-034.

Black:1957:ABLc

[291] Edgar C. Black. Alterations in the blood level of lactic acid in certain salmonoid fishes following muscular activity: III. Sockeye salmon, Oncorhynchus nerka. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6): 807–814, June 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f57-035.

Hoar:1957:RJP

[292] W. S. Hoar, M. H. A. Keenleyside, and R. G. Goodall. Reactions of juvenile Pacific salmon to light. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6):815–830, June 1957. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f57-036.

Templeman:1957:NFC

[293] Wilfred Templeman, H. J. Squires, and A. M. Fleming. Nematodes in the fillets of cod and other fishes in Newfoundland and neighbouring areas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6):831–897, June 1957.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f57-037.

Steele:1957:RSM

[294] D. H. Steele. The redfish (Sebastes marinus L.) in the Western Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6): 899–924, June 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f57-038.

Dyer:1957:FSI

[295] W. J. Dyer, Doris I. Fraser, and W. A. MacCallum. Further study of the influence of short storage periods, 3 days to 2 Weeks at 15^{o}f., on the quality of frozen cod stored at 0^{o}F. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6):925–929, June 1957. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f57-039.

Bumpus:1957:NTS

[296] Dean F. Bumpus, Joseph Chase, C. Godfrey Day, David H. Frantz, Jr., David D. Ketchum, and Robert G. Walden. A new technique for studying non-tidal drift with results of experiments off gay head, mass., and in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6): 931–944, June 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f57-040.

Barber:1957:EPW

[297] F. G. Barber. The effect of the prevailing winds on the inshore water masses of the Hecate Strait Region, B.C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6):945–952, June 1957. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f57-041.

Alverson:1957:RTE

[298] Dayton L. Alverson and Bruce M. Chatwin. Results from tagging experiments on a spawning stock of petrale sole, Eopsetta jordani (Lockington). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6):953–974, June 1957. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f57-042.

Scott:1957:VIL

[299] D. M. Scott and W. R. Martin. Variation in the incidence of larval nematodes in Atlantic cod fillets along the Southern Canadian Mainland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6):975–996, June 1957. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f57-043.

Anonymous:1957:CRP

[300] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6): 997–998, June 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f57-044.

Anonymous:1957:IVJ

[301] Anonymous. Index to volume 14 of the Journal of the Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 14(6): 999–1008, June 1957. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f57-045.

Scott:1958:IPA

[302] D. M. Scott and H. D. Fisher. Incidence of a parasitic ascarid, Porrocaecum decipiens, in the common porpoise, Phocoena phocoena, from the Lower Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(1): 1–4, January 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f58-001.

Pike:1958:ADN

[303] Gordon C. Pike and Brian E. Maxwell. The abundance and distribution of the northern sea lion (Eumetopias jubata) on the coast of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(1):5–17, January 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f58-002.

DeLury:1958:EPS

[304] D. B. DeLury. The estimation of population size by a marking and recapture procedure. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(1):

19–25, January 1958. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f58-003.

Miller:1958:RCM

[305] Richard B. Miller. The Rôle of competition in the mortality of hatchery trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(1):27–45, January 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f58-004.

Vladykov:1958:RLA

[306] Vadim D. Vladykov and W. I. Follett. Redescription of Lampetra ayresii (Günther) of Western North America, a species of lamprey (Petromyzontidae) distinct from Lampetra fluviatilis (Linnaeus) of Europe. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(1):47–77, January 1958. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f58-005.

Trites:1958:CSS

[307] R. W. Trites and R. E. Banks. Circulation on the Scotian Shelf as indicated by drift bottles. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(1): 79–89, January 1958. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f58-006.

Tabata:1958:SWI

[308] Susumu Tabata and Robin J. LeBrasseur. Sea water intrusion into the Fraser River and its relation to the incidence of shipworms in Steveston Cannery Basin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(1):91– 113, January 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f58-007.

McLellan:1958:ECB

[309] H. J. McLellan. Energy considerations in the Bay of Fundy system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(2):115–134, February 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f58-008.

Dingle:1958:PFM

[310] J. R. Dingle. Proteins in fish muscle. 11. On the extraction of myosin from cod muscle with potassium iodide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(2):135–153, February 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-009.

Dunbar:1958:POR

[311] M. J. Dunbar. Physical oceanographic results of the “Calanus” Expeditions in Ungava Bay, Frobisher Bay, Cumberland Sound, Hudson Strait and Northern Hudson Bay, 1949–1955. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(2):155–201, February 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-010.

Lawler:1958:LSH

[312] G. H. Lawler and N. H. F. Watson. Limnological studies of Heming Lake, Manitoba, and two adjacent lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(2):203–218, February 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-011.

McLaren:1958:SAG

[313] I. A. McLaren. Some aspects of growth and reproduction of the bearded seal, Erignathus barbatus (Erxleben). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(2):219–227, February 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-012.

Alderdice:1958:SET

[314] D. F. Alderdice, W. P. Wickett, and J. R. Brett. Some effects of temporary exposure to low dissolved oxygen levels on Pacific salmon eggs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(2):229–250, February 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f58-013.

Hoar:1958:RLC

[315] William S. Hoar. Rapid learning of a constant course by travelling schools of juvenile Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(2):251–274, February 1958. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f58-014.

Pitt:1958:DSR

[316] T. K. Pitt. Distribution, spawning and racial studies of the capelin, Mallotus villosus (Müller), in the offshore Newfoundland area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(3):275–293, March 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-015.

Pitt:1958:AGC

[317] T. K. Pitt. Age and growth of the capelin, Mallotus villosus (Müller), from Newfoundland and Grand Bank areas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(3):295–311, March 1958. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f58-016.

Nagasaki:1958:FPH

[318] Fuzuko Nagasaki. The fecundity of Pacific herring (Clupea pallasi) in British Columbia coastal waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(3):313–330, March 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-017.

Godfrey:1958:CSS

[319] Harold Godfrey. A comparison of sockeye salmon catches at Rivers Inlet and Skeena River, B.C., with particular reference to age at maturity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(3):331–354, March 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f58-018.

Templeman:1958:VFL

[320] Wilfred Templeman and V. M. Hodder. Variation with fish length, sex, stage of sexual maturity, and season in the appearance and volume of the drumming muscles of the swim-bladder in the haddock, Melanogrammus aeglefinus (L.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(3): 355–390, March 1958. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f58-019.

Hoar:1958:EMB

[321] William S. Hoar. The evolution of migratory behaviour among juvenile salmon of the genus Oncorhynchus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(3):391–428, March 1958. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f58-020.

Tabata:1958:HBW

[322] Susumu Tabata. Heat budget of the water in the vicinity of triple Island, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(3): 429–451, March 1958. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f58-021.

Strickland:1958:SRP

[323] J. D. H. Strickland. Solar radiation penetrating the ocean. A review of requirements, data and methods of measurement, with particular reference to photosynthetic productivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(3):453–493, March 1958. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f58-022.

Scott:1958:IAP

[324] D. M. Scott and H. D. Fisher. Incidence of the ascarid Porrocaecum decipiens in the stomachs of three species of seals along the Southern Canadian Atlantic Mainland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):495–516, April 1958. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f58-023.

Templeman:1958:RHL

[325] Wilfred Templeman and A. M. Fleming. Round haddock landings in Newfoundland: Quantity and quality relative to gutted haddock. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):517–527, April 1958. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f58-024.

Chew:1958:PSF

[326] Kenneth K. Chew and Ronald Eisler. A preliminary study of the feeding habits of the Japanese oyster drill, Ocinebra japonica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):529–535, April 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-025.

Power:1958:EVL

[327] H. E. Power. The effect of various lighting conditions on the efficiency of “candling” cod fillets for detection of parasites. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):537–542, April 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-026.

Legendre:1958:ADS

[328] R. Legendre. Artificial drying of salt fish by thermocouple control. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):543–554, April 1958. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f58-027.

Cardin:1958:CFA

[329] A. Cardin, M. A. Bordeleau, and A. Laframboise. The composition of the fatty acids liberated from the phospholipids in the preparation of salted codfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4): 555–558, April 1958. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f58-028.

Beak:1958:TFT

[330] T. W. Beak. Toleration of fish to toxic pollution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):559–572, April 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-029.

Black:1958:HLF

[331] Edgar C. Black. Hyperactivity as a lethal factor in fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):573–586, April 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-030.

Brett:1958:ETC

[332] J. R. Brett, M. Hollands, and D. F. Alderdice. The effect of temperature on the cruising speed of young sockeye and Coho salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):587–605, April 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-031.

Ferguson:1958:PTF

[333] R. G. Ferguson. The preferred temperature of fish and their midsummer distribution in temperate lakes and streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):607–624, April 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-032.

Bell:1958:CTC

[334] F. Heward Bell and Alonzo T. Pruter. Climatic temperature changes and commercial yields of some marine fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):625–683, April 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-033.

Lindahl:1958:PRL

[335] Per Eric Lindahl and Arne Marcström. On the preference of roaches (Leuciscus rutilus) for trinitrophenol, studied with the fluviarium technique. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):685–694, April 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f58-034.

Templeman:1958:RFR

[336] Wilfred Templeman and E. J. Sandeman. Red flesh in redfish, Sebastes marinus (L.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4): 695–700, April 1958. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f58-035.

Castell:1958:GFQa

[337] C. H. Castell, Maxine F. Greenough, R. S. Rodgers, and A. S. MacFablane. Grading fish for quality. 1. Trimethylamine values of fillets cut from graded fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4): 701–716, April 1958. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f58-036.

Hoogland:1958:GFQ

[338] P. L. Hoogland. Grading fish for quality. 2. Statistical analysis of the results of experiments regarding grades and trimethylamine values. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):717–728, April 1958. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f58-037.

Castell:1958:GFQb

[339] C. H. Castell and Maxine F. Greenough. Grading fish for quality. 3. Grading of recently cut fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):729–748, April 1958. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f58-038.

Paloheimo:1958:MEN

[340] J. E. Paloheimo. A method of estimating natural and fishing mortalities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4):749–758, April 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f58-039.

Anonymous:1958:EFa

[341] Anonymous. Editors’ foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(4), April 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f58-040.

Needler:1958:FRB

[342] A. W. Needler. Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):759–777, May 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-041.

Clemens:1958:RD

[343] W. A. Clemens. Reminiscences of a director. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):779–796, May 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-042.

Alderdice:1958:NRD

[344] D. F. Alderdice and W. P. Wickett. A note on the response of developing chum salmon eggs to free carbon dioxide in solution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):797–799, May 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-043.

Berkeley:1958:PWC

[345] E. Berkeley and C. Berkeley. Polychaeta of the Western Canadian Arctic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):801–804, May 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f58-044.

Brett:1958:RCY

[346] J. R. Brett and D. F. Alderdice. The resistance of cultured young chum and sockeye salmon to temperatures below 0 ^{o}C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):805–813, May 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-045.

Brett:1958:SOS

[347] J. R. Brett and M. A. Ali. Some observations on the structure and photomechanical responses of the Pacific salmon retina. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):815–829, May 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-046.

Chatwin:1958:MRE

[348] Bruce M. Chatwin. Mortality rates and estimates of theoretical yield in relation to minimum commercial size of lingcod (Ophiodon elongatus) from the Strait of Georgia, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):831–849, May 1958. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f58-047.

Dodimead:1958:PRD

[349] A. J. Dodimead and H. J. Hollister. Progress report of drift bottle releases in the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):851–865, May 1958. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f58-048.

Foskett:1958:RIS

[350] D. R. Foskett. The Rivers Inlet sockeye salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):867–889, May 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-049.

Godfrey:1958:CIR

[351] H. Godfrey. Comparisons of the index of return for several Stocks of British Columbia salmon to study variations in survival. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):891–908, May 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-050.

Hourston:1958:PSJ

[352] Alan S. Hourston. Population studies on juvenile herring in Barkley Sound, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):909–960, May 1958. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f58-051.

Johnson:1958:DDY

[353] W. E. Johnson. Density and distribution of young sockeye salmon (Oncorhynchus nerka) throughout a multibasin lake system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):961–982, May 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-052.

Margolis:1958:OJC

[354] L. Margolis. The occurrence of juvenile Corynosoma (Acanthocephala) in Pacific salmon (Oncorhynchus spp.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):983–990, May 1958. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f58-053.

Ricker:1958:MSY

[355] W. E. Ricker. Maximum sustained yields from fluctuating environments and mixed stocks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5): 991–1006, May 1958. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f58-054.

Shepard:1958:SSS

[356] M. P. Shepard and F. C. Withler. Spawning stock size and resultant production for skeena sockeye. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):1007–1025, May 1958. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f58-055.

Terhune:1958:MVG

[357] L. D. B. Terhune. The mark VI groundwater standpipe for measuring seepage through salmon spawning gravel. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):1027–1063, May 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-056.

Waldichuk:1958:DBO

[358] Michael Waldichuk. Drift bottle observations in the Strait of Georgia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):1065–1102, May 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f58-057.

Wickett:1958:RCE

[359] W. P. Wickett. Review of certain environmental factors affecting the production of pink and chum salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(5):1103–1126, May 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-058.

Anonymous:1958:EFb

[360] Anonymous. Editors’ foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6), June 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f58-059.

Hart:1958:FRB

[361] J. L. Hart. Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1127–1161, June 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-060.

Bailey:1958:DFS

[362] W. B. Bailey. On the dominant flow in the Strait of Belle Isle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1163–1174, June 1958. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f58-061.

Campbell:1958:DFB

[363] N. J. Campbell and A. E. Collin. The discoloration of Foxe Basin ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1175–1188, June 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f58-062.

Dickie:1958:EHT

[364] L. M. Dickie. Effects of high temperature on survival of the giant scallop. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1189–1211, June 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f58-063.

Huntsman:1958:SS

[365] A. G. Huntsman. Shubenacadie salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1213–1218, June 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-064.

Jermolajev:1958:ZIB

[366] E. G. Jermolajev. Zooplankton of the inner Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1219–1228, June 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-065.

Kohler:1958:VOA

[367] A. C. Kohler. The validity of otolith age determinations for haddock (Melanogrammus aeglefinus L.) from the lockeport, N.S., area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1229–1238, June 1958. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f58-066.

Kohler:1958:HSO

[368] A. C. Kohler and J. R. Clark. Haddock scale–otolith comparisons. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1239–1246, June 1958. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f58-067.

Lauzier:1958:DWL

[369] L. M. Lauzier and R. W. Trites. The deep waters in the Laurentian Channel. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1247– 1257, June 1958. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f58-068.

Leim:1958:FSH

[370] A. H. Leim. Fatness of small herring in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1259–1267, June 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-069.

McCracken:1958:BFC

[371] F. D. McCracken. On the biology and fishery of the Canadian Atlantic halibut, Hippoglossus hippoglossus L. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1269–1311, June 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-070.

McKenzie:1958:AGS

[372] R. A. McKenzie. Age and growth of smelt, Osmerus mordax (Mitchill), of the Miramichi River, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1313–1327, June 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-071.

McKenzie:1958:HMP

[373] R. A. McKenzie and B. E. Skud. Herring migrations in the Passamaquoddy region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):

1329–1343, June 1958. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f58-072.

McLeese:1958:ACL

[374] D. W. McLeese and D. G. Wilder. The activity and catchability of the lobster (Homarus americanus) in relation to temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1345–1354, June 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-073.

Medcof:1958:TCG

[375] J. C. Medcof and L. W. Thurber. Trial control of the greater clam drill (Lunatia heros) by manual collection. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1355–1369, June 1958. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f58-074.

Paloheimo:1958:DNF

[376] J. E. Paloheimo. Determination of natural and fishing mortalities of cod and haddock from analysis of tag records off Western Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1371–1381, June 1958. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f58-075.

Powles:1958:SRF

[377] P. M. Powles. Studies of reproduction and feeding of Atlantic cod (Gadus callarias L.) in the Southwestern Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1383–1402, June 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-076.

Smith:1958:MBT

[378] M. W. Smith and J. W. Saunders. Movements of brook trout, Salvelinus fontinalis (Mitchill), between and within fresh and salt water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1403–1449, June 1958. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f58-077.

Tibbo:1958:ODL

[379] S. N. Tibbo, J. E. Henri Legaré, Leslie W. Scattergood, and R. F. Temple. On the occurrence and distribution of larval herring (Clupea harengus L.) in the Bay of Fundy and the Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 15(6):1451–1469, June 1958. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f58-078.

Visweswariah:1959:ECA

[380] K. Visweswariah, M. N. Moorjani, D. S. Bhatia, and V. Subrahmanyan. Effect of chlortetracycline (aureomycin) on the keeping quality of freshwater fish under tropical conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):1–5, January 1959. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f59-001.

Young:1959:CCS

[381] E. Gordon Young, D. G. Smith, and W. M. Langille. The chemical composition of sea water in the vicinity of the Atlantic Provinces of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):7–12, January 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f59-002.

Castell:1959:APFa

[382] C. H. Castell, Maxine F. Greenough, and Jacqueline Dale. The action of Pseudomonas on fish muscle: 3. Identification of organisms producing fruity and oniony odours. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):13–19, January 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-003.

Castell:1959:APFb

[383] C. H. Castell and Maxine F. Greenough. The action of Pseudomonas on fish muscle: 4. Relation between substrate composition and the development of odours by Pseudomonas fragi. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):21–31, January 1959. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f59-004.

Ellis:1959:PFM

[384] D. G. Ellis and P. M. Winchester. Proteins in fish muscle. 12. Ultracentrifuge studies on post-rigor extracts of structural protein. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):33–41, January 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-005.

Dyer:1959:PFM

[385] W. J. Dyer and Doris I. Fraser. Proteins in fish muscle. 13. Lipid hydrolysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):43–52, January 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f59-006.

Dingle:1959:PFMa

[386] J. R. Dingle and P. H. Odense. Proteins in fish muscle. 14. Cod tropomyosin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):53–62, January 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f59-007.

Dambergs:1959:EFM

[387] N. Dambergs. Extractives of fish muscle. 2. Solvent–water ratio in extraction of fat and water-solubles. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):63–71, January 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-008.

Bishop:1959:EEM

[388] Yvonne M. M. Bishop. Errors in estimates of mortality obtained from virtual populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1): 73–90, January 1959. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f59-009.

Anas:1959:TYO

[389] Raymond E. Anas. Three-year-old pink salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):91–94, January 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-010.

Parker:1959:MFM

[390] Robert R. Parker and Edgar C. Black. Muscular fatigue and mortality in troll-caught Chinook salmon (Oncorhynchus tshawytscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):95–106, January 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-011.

Hewson:1959:SYS

[391] L. C. Hewson. A seven-year study of the fishery for lake whitefish, Coregonus clupeaformis, on Lake Winnipeg. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):107–120, January 1959. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f59-012.

Canagaratnam:1959:GFD

[392] P. Canagaratnam. Growth of fishes in different salinities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):121–130, January 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-013.

Hewson:1959:SSW

[393] L. C. Hewson. A study of six winter seasons of commercial fishing on Lake Winnipeg, 1950–1955. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(1):131–145, January 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-014.

Hallam:1959:HAF

[394] J. C. Hallam. Habitat and associated fauna of four species of fish in Ontario streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(2): 147–173, February 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-015.

Basu:1959:ARF

[395] Satyendra Prasanna Basu. Active respiration of fish in relation to ambient concentrations of oxygen and carbon dioxide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(2):175–212, February 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-016.

Scott:1959:INF

[396] D. M. Scott and W. R. Martin. The incidence of nematodes in the fillets of small cod from Lockeport, Nova Scotia, and the Southwestern Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(2): 213–221, February 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-017.

Castell:1959:SFV

[397] C. H. Castell, Jaqueline Dale, and Maxine F. Greenough. Spoilage of fish in the vessels at sea: 6. Variations in the landed quality of trawler-caught Atlantic cod and haddock during a period of 13 months. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(2):223–233, February 1959. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f59-018.

Idler:1959:BSSa

[398] D. R. Idler and I. Bitners. Biochemical studies on sockeye salmon during spawning migration. V. Cholesterol, fat, protein and water in the body of the standard fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(2): 235–241, February 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-019.

Dingle:1959:PFMb

[399] J. R. Dingle. Proteins in fish muscle. 15. Note on the preparation of actin from cod muscle with potassium iodide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(2):243–246, February 1959. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f59-020.

Crossman:1959:DMP

[400] E. J. Crossman. Distribution and movements of a predator, the rainbow trout, and its prey, the redside shiner, in Paul Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3):247–267, March 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f59-021.

Crossman:1959:PPI

[401] E. J. Crossman. A predator–prey interaction in freshwater fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3):269–281, March 1959. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f59-022.

Hourston:1959:ESA

[402] Alan S. Hourston. Effects of some aspects of environment on the distribution of juvenile herring in Barkley Sound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3):283–308, March 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-023.

Hourston:1959:RJH

[403] Alan S. Hourston. The relationship of the juvenile herring stocks in Barkley Sound to the major adult herring populations in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3):309– 320, March 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f59-024.

Miller:1959:DGR

[404] R. B. Miller, A. C. Sinclair, and P. W. Hochachka. Diet, glycogen reserves and resistance to fatigue in hatchery rainbow trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3):321–328, March 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-025.

Godfrey:1959:VAAa

[405] Harold Godfrey. Variations in annual average weights of British Columbia pink salmon, 1944–1958. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3):329–337, March 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-026.

Olsen:1959:MSH

[406] Steinar Olsen. Mesh selection in herring gill nets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3):339–349, March 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-027.

Sund:1959:DCG

[407] Paul N. Sund. The distribution of Chaetognatha in the Gulf of Alaska in 1954 and 1956. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3): 351–361, March 1959. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f59-028.

Briggs:1959:MOL

[408] John C. Briggs and Nobuo Egami. The medaka (Oryzias latipes). A commentary and a bibliography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3):363–380, March 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-029.

Hoy:1959:PCB

[409] Joe Hoy. A portable conductivity bridge for field use. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3):381–382, March 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-030.

Neave:1959:RFW

[410] Ferris Neave. Records of fishes from waters off the British Columbia Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3):383–384, March 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f59-031.

Anonymous:1959:CRPa

[411] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(3): 385–386, March 1959. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f59-032.

Anonymous:1959:O

[412] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):387–390, April 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-033.

Black:1959:CPC

[413] Edgar C. Black, Wing gay Chiu, Francis D. Forbes, and Arthur Hanslip. Changes in pH, carbonate and lactate of the blood of yearling Kamloops trout, Salmo gairdneri, during and following severe muscular activity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):391–402, April 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f59-034.

Grainger:1959:SLF

[414] E. H. Grainger and J. G. Hunter. Station list of the 1955–58 field investigations of the Arctic Unit of the Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):403–420, April 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f59-035.

Creelman:1959:BSS

[415] Vera M. Creelman and Neil Tomlinson. Biochemical studies on sockeye salmon during spawning migration: VI. Ribonucleic acid and deoxyribonucleic acid. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4): 421–428, April 1959. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f59-036.

Parker:1959:FMT

[416] R. R. Parker, E. C. Black, and P. A. Larkin. Fatigue and mortality in trollcaught Pacific salmon (Oncorhynchus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):429–448, April 1959. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f59-037.

Bligh:1959:ORF

[417] E. G. Bligh and W. J. Dyer. Orange-red flesh in cod and haddock. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):449–452, April 1959. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f59-038.

Grainger:1959:AOC

[418] E. H. Grainger. The annual oceanographic cycle at Igloolik in the Canadian Arctic. 1. The zooplankton and physical and chemical observations.

Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):453–501, April 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f59-039.

Leim:1959:RUU

[419] A. H. Leim and L. R. Day. Records of uncommon and unusual fishes from Eastern Canadian Waters, 1950–1958. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):503–514, April 1959. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f59-040.

Berg:1959:VHR

[420] L. S. Berg. Vernal and hiemal races among anadromous fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):515–537, April 1959. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f59-041.

Castell:1959:FAN

[421] C. H. Castell and Maxine F. Greenough. Factors affecting the nitrite content of treated fish fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):539–552, April 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-042.

Godfrey:1959:VAAb

[422] H. Godfrey. Variation in the annual average weight of chum salmon caught in British Columbia waters, 1946 to 1958. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):553–554, April 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-043.

Shewan:1959:CCS

[423] J. M. Shewan and A. S. C. Ehrenberg. Comparison of certain Scottish and Canadian experiments in respect of grading fish for quality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):555–557, April 1959. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f59-044.

Idler:1959:BSSb

[424] D. R. Idler and H. Tsuyuki. Biochemical studies on soekeye salmon during spawning migration: VIII. Androgen content of testes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):559–560, April 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-045.

McBride:1959:SCF

[425] J. R. McBride, R. A. MacLeod, and D. R. Idler. Studies on the conversion of fish stickwater to solubles: IV. Effect of enzyme treatment on the viscosity of whale solubles. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(4):561–563, April 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-046.

Bennett:1959:SOF

[426] E. B. Bennett. Some oceanographic features of the Northeast Pacific Ocean during August 1955. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(5):565–633, May 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-047.

Pickard:1959:CMK

[427] G. L. Pickard and Keith Rodgers. Current measurements in Knight Inlet, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(5): 635–678, May 1959. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f59-048.

McBride:1959:PAP

[428] J. R. McBride, R. A. MacLeod, and D. R. Idler. Proximate analysis of Pacific herring (Clupea pallasii) and an evaluation of Tester’s “Fat Factor”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(5):679–684, May 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f59-049.

MacLeod:1959:CFS

[429] Robert A. MacLeod. Condensed fish solubles. A review of its preparation and properties. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(5):

685–694, May 1959. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f59-050.

Lindsey:1959:HRT

[430] C. C. Lindsey, T. G. Northcote, and G. F. Hartman. Homing of rainbow trout to inlet and outlet spawning streams at Loon Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(5):695– 719, May 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f59-051.

Parker:1959:CGF

[431] Robert R. Parker and Peter A. Larkin. A concept of growth in fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(5):721–745, May 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f59-052.

Bilinski:1959:ESC

[432] E. Bilinski and H. Fougère. The effect of sodium chloride on proteolysis and on the fate of amino acids present in the muscle of codfish (Gadus callarias). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(5):747– 754, May 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f59-053.

Wood:1959:DLE

[433] J. D. Wood. Distribution of a lipase enzyme in lingcod fillets and the effect of low temperature storage on its activity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(5):755–757, May 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-054.

Barraclough:1959:FRN

[434] W. E. Barraclough. The first record of a northern blennioid fish, Plectobranchus evides Gilbert (family Stichaeidae), in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(5):759–760, May 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f59-055.

Butler:1959:RAC

[435] T. H. Butler. A record of the anomuran crustacean Emerita analoga (Stimpson) from British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(5):761, May 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-056.

Templeman:1959:VCP

[436] Wilfred Templeman and E. J. Sandeman. Variations in caudal pigmentation in late-stage pre-extrusion larvae from marinus- and mentellatype female redfish from the Newfoundland area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6):763–789, June 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-057.

Moyer:1959:KQP

[437] R. H. Moyer, B. A. Southcott, E. G. Baker, and H. L. A. Tarr. Keeping quality of Pacific Coast dogfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6):791–794, June 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-058.

LeBrasseur:1959:SLB

[438] R. J. LeBrasseur. Sagitta lyra, a biological indicator species in the subarctic waters of the eastern Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6):795–805, June 1959. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f59-059.

McKenzie:1959:MFF

[439] R. A. McKenzie. Marine and freshwater fishes of the Miramichi River and Estuary, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6):807–833, June 1959. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f59-060.

Hunter:1959:SPP

[440] J. G. Hunter. Survival and production of pink and chum salmon in a coastal stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6): 835–886, June 1959. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f59-061.

Smith:1959:PED

[441] M. W. Smith. Phosphorus enrichment of drainage waters from farm lands. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6):887–895, June 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f59-062.

Ricker:1959:AOC

[442] W. E. Ricker. Additional observations concerning residual sockeye and Kokanee (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6):897–902, June 1959. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f59-063.

Qadri:1959:SMD

[443] S. U. Qadri. Some morphological differences between the subspecies of cutthroat trout, Salmo clarkii clarkii and Salmo clarkii lewisi, in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6):903– 922, June 1959. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f59-064.

Neumann:1959:MDA

[444] J. Neumann. Maximum depth and average depth of lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6):923–927, June 1959. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f59-065.

Anonymous:1959:CRPb

[445] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6): 929–931, June 1959. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f59-066.

Anonymous:1959:IVJ

[446] Anonymous. Index to volume 16 of the Journal of the Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 16(6): 933–944, June 1959. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f59-067.